### v1

# Distance of Quadratic Algebraic Numbers from the Middle-Third Cantor Set

Abstract
We study the distance from quadratic irrational numbers to the middle-third Cantor set $C$. Mahler asked whether $C$ contains any irrational algebraic numbers; this remains open even for quadratic irrationals. Rather than assuming an answer to this problem, we obtain uniform lower bounds for the distance from a quadratic irrational $\alpha$ to $C$ in terms of the height $H$ of the minimal polynomial of $\alpha$.
We encode $\alpha$ by its orbit under the map $x\mapsto 3x\bmod 1$ and define the exit time $\operatorname{exit}(\alpha)$ as the first iterate that enters the middle interval $[1/3,2/3]$. Our main unconditional result is a quadratic exit bound $\operatorname{exit}(\alpha)\leq A(\log_{3}H)^{2}+B$ for absolute constants $A,B>0$, valid for all quadratic irrationals whose orbit stays a fixed small distance away from the coarse Cantor boundaries. As a consequence we obtain a distance lower bound $\operatorname{dist}(\alpha,C)\geq H^{-\kappa\log H}$ for some constant $\kappa>0$.
On the dynamical side we classify orbits by an $L/M/R$ coding and prove that the total number of visits to the right interval $[2/3,1)$ is $O(\log H)$. A finite case analysis on a bounded portion of the orbit is reduced to checking a finite list of explicit affine inequalities on subintervals of $[0,1]$, which we verify with short computer scripts; all Diophantine and dynamical estimates are proved by hand.
MSC 2020: 11J04 (primary), 28A80
Keywords: Cantor set, algebraic numbers, Mahler’s problem, Diophantine approximation, quadratic irrationals
1 Introduction
Let $\mathcal{C}\subset[0,1]$ denote the standard middle-third Cantor set. It is a classical object of fractal geometry, of Hausdorff dimension $\log 2/\log 3$, arising by repeatedly removing the open middle third from each closed interval. In base $3$ it can be described as
$\mathcal{C}=\Big\{\theta\in[0,1]:\theta=0.\varepsilon_{1}\varepsilon_{2}\cdots\text{ in base }3\text{ with }\varepsilon_{n}\in\{0,2\}\ \forall n\Big\}.$
It is a long-standing open problem, going back to Mahler, to understand the algebraic points of $\mathcal{C}$. In his 1984 paper on Diophantine approximation on Cantor-like sets Mahler asked whether $\mathcal{C}$ contains any irrational algebraic numbers at all [3]. This remains wide open: in particular, it is not known whether a quadratic irrational can lie in $\mathcal{C}$. Levesley–Salp–Velani explicitly note that, to the best of current knowledge, it is not even known whether or not quadratic irrationals avoid the middle-third Cantor set [4, Remark 1.7]. Our results therefore do not presuppose that algebraic points are excluded from $\mathcal{C}$; rather, we obtain quantitative lower bounds for the distance from $\mathcal{C}$ that hold uniformly for all quadratic irrationals whose orbits stay a fixed distance away from the coarse Cantor boundaries.
1.1 Exit time and dynamical formulation
Let $\alpha\in\mathbb{R}$ and consider the orbit
$\theta_{N}\coloneqq\{3^{N}\alpha\}\in[0,1],\qquad N=0,1,2,\dots.$
We partition $[0,1]$ into three intervals
$L\coloneqq[0,1/3),\quad M\coloneqq[1/3,2/3),\quad R\coloneqq[2/3,1).$
The Cantor set $\mathcal{C}$ is the set of all $\theta_{0}$ whose orbit under the map $\theta\mapsto\{3\theta\}$ never enters $M$. Thus for $\theta_{0}\notin\mathcal{C}$ the first time we enter $M$ is a natural dynamical measure of how far $\theta_{0}$ is from $\mathcal{C}$.
Definition 1.1.
Let $\alpha\in\mathbb{R}$. We define the exit time
$\mathrm{exit}(\alpha)\coloneqq\inf\{N\geq 1:\theta_{N-1}\in M\}\in\mathbb{N}\cup\{\infty\}$
with the convention that $\mathrm{exit}(\alpha)=\infty$ if the orbit never enters $M$. When $\mathrm{exit}(\alpha)<\infty$ we say that $\alpha$ exits by time $N=\mathrm{exit}(\alpha)$.
Thus $\mathrm{exit}(\alpha)=1$ if and only if $\theta_{0}\in M$, and more generally $\mathrm{exit}(\alpha)=N$ if and only if $\theta_{N-1}\in M$ while $\theta_{0},\dots,\theta_{N-2}\notin M$.
1.2 Height and coarse boundary distance
We restrict attention to quadratic irrational numbers.
Definition 1.2.
Let $\alpha$ be a quadratic irrational. We write $f_{\alpha}(x)=ax^{2}+bx+c\in\mathbb{Z}[x]$ for its primitive minimal polynomial (with $\gcd(a,b,c)=1$ and $a>0$), and we define the height $H(\alpha)$ to be
$H(\alpha)\coloneqq\max\{|a|,|b|,|c|\}.$
We will often abbreviate $H(\alpha)$ to $H$ when $\alpha$ is fixed.
The middle-third Cantor set has a natural hierarchical structure: at level $1$ we have boundary points $\{0,1/3,2/3,1\}$, at level $2$ we have $\{k/9:0\leq k\leq 9\}$, and so on. In this paper it is convenient to measure distance to $\mathcal{C}$ at level $2$.
Definition 1.3.
For $\theta\in[0,1]$ we define the coarse boundary distance
$\delta_{\mathcal{C}}(\theta)\coloneqq\min_{0\leq k\leq 9}|\theta-k/9|.$
For a quadratic irrational $\alpha$ we set $\delta_{\mathcal{C}}(\alpha)\coloneqq\delta_{\mathcal{C}}(\{\alpha\})$.
Note that $\delta_{\mathcal{C}}$ is not the true distance to $\mathcal{C}$, but a coarse proxy that measures how close we are to the level-$2$ Cantor boundaries. The regime $\delta_{\mathcal{C}}(\alpha)<0.02$ is obstructed by rational Cantor points and is not treated in this paper; see Section 6 for a brief discussion.
The specific numerical thresholds used below,
$\delta_{\mathcal{C}}(\theta)\geq 0.02\quad\text{and}\quad\delta_{\mathcal{C}}(\theta)\geq 0.04,$
have no intrinsic arithmetic significance. Any fixed choice of constants $0<c_{1}<c_{2}<1/9$ would serve equally well for our arguments, after adjusting the implicit constants in the final bounds. We fix $0.02$ and $0.04$ simply because they are comfortably smaller than the level-$2$ mesh size $1/9$ and interact well with the finite interval partitions used in the proofs of Theorem 3.1 and Proposition 5.1 (see Appendix A). In particular, all the dynamical and Diophantine estimates are uniform in the sense that the choice of $0.02$ could be replaced by any other small absolute constant.
1.3 Main results
Our main object of study is the relationship between $\mathrm{exit}(\alpha)$, the height $H(\alpha)$ and the distance from $\alpha$ to $\mathcal{C}$.
Theorem 1.4 (Generic exit bound).
Let $\alpha$ be a quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ and height $H(\alpha)=H$. Suppose that $3\nmid a$ and $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then
$\mathrm{exit}(\alpha)\leq\log_{3}H+1.$
This theorem corresponds to the “generic” case in which the leading coefficient is not divisible by $3$. In this regime the 3-adic structure is simple and a direct computation suffices to bound the number of steps before the orbit enters $M$.
Our next theorem treats the general quadratic case without any restriction on the 3-adic valuation of the coefficients, still under the coarse boundary condition $\delta_{\mathcal{C}}(\alpha)\geq 0.02$.
Theorem 1.5 (General quadratic exit bound).
There exist absolute constants $A,B>0$ such that the following holds. Let $\alpha$ be any quadratic irrational with minimal polynomial of height $H(\alpha)=H$ and with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then
$\mathrm{exit}(\alpha)\leq A(\log_{3}H)^{2}+B.$
As a consequence we obtain a general distance lower bound.
Corollary 1.6 (General distance bound).
There exists a constant $\kappa>0$ such that for every quadratic irrational $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and height $H(\alpha)=H$ we have
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}.$
Here $\mathrm{dist}(\alpha,\mathcal{C})$ denotes the Euclidean distance from $\alpha$ to the Cantor set.
In particular, Theorem 1.5 and Corollary 1.6 give the first unconditional general lower bound $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}$ for quadratic irrationals (away from the small-$\delta_{\mathcal{C}}$ obstructions).
We also identify a special “baby resonant” regime in which the 3-adic structure of the quadratic polynomial has a particularly simple form (roughly, $v_{3}(a)=1$ and $v_{3}(c)=0$). In this setting we obtain a conditional logarithmic exit bound under a natural uniform bound on the number of R-blocks; see Section 4. While the full unconditional logarithmic behaviour remains conjectural, the baby resonant analysis plays an important role in clarifying the dynamical mechanisms that govern exit times.
1.4 Fractal dimension and Cantor background
For completeness we recall that the middle-third Cantor set $\mathcal{C}$ has Hausdorff dimension
$\dim_{H}(\mathcal{C})=\frac{\log 2}{\log 3}.$
Our dynamical formulation can be viewed as studying how the orbit of a point under the expanding map $\theta\mapsto\{3\theta\}$ interacts with the “holes” defining $\mathcal{C}$. The scale $3^{-N}$ appearing in the definition of $\mathrm{exit}(\alpha)$ corresponds to the level-$N$ intervals in the standard Cantor construction, and the distance estimates in Corollary 1.6 should be compared with the natural Diophantine scale $H^{-2}$ for quadratic irrationals.
1.5 Nature of the computer-assisted steps
A few remarks about the computer-assisted parts of the proofs are in order. The only places where we invoke a computer are in the proofs of Theorem 3.1 and Proposition 5.1. In both cases the role of the code is to verify a finite list of explicit affine inequalities arising from a partition of the unit interval into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is fixed. Conceptually, these steps are standard “finite case analyses” in the spirit of a computer-aided dissection of finitely many dynamical regimes; all the Diophantine and dynamical estimates are proved by hand.
For transparency, Appendix A explains in detail how the relevant finite collections of intervals are constructed and what statements are actually checked, and provides references to the short Python scripts that carry out the verification. The scripts are elementary and may be inspected or reimplemented independently if desired. No external computer algebra system or deep symbolic computation is used.
Acknowledgements
[Omitted here.]
2 Dynamical setup and basic lemmas
In this section we record the basic dynamical facts about the map $\theta\mapsto\{3\theta\}$ and its interaction with the partition $L\cup M\cup R$.
2.1 Digit index convention
We recall Definition 1.1 and the digit-index convention. For completeness we record the following simple observation.
Remark 2.1.
By definition we have $\mathrm{exit}(\alpha)=N$ if and only if $\theta_{N-1}\in M$ and $\theta_{j}\notin M$ for $0\leq j\leq N-2$. Thus bounds on the first time $N_{0}$ with $\theta_{N_{0}}\in M$ translate into exit-time bounds via $\mathrm{exit}(\alpha)=N_{0}+1$.
We will systematically work in the $\theta_{N}$-index convention and translate back to $\mathrm{exit}(\alpha)$ only at the end of proofs.
2.2 Peeling near the endpoints
We recall that
$L=[0,1/3),\quad M=[1/3,2/3),\quad R=[2/3,1).$
On each of these intervals the map $\theta\mapsto\{3\theta\}$ is affine. In particular, if $\theta_{0}\in L$ then $\theta_{1}=3\theta_{0}\in[0,1)$, if $\theta_{0}\in M$ then $\theta_{1}=3\theta_{0}-1$, and if $\theta_{0}\in R$ then $\theta_{1}=3\theta_{0}-2$.
We will frequently use the following simple but useful observation.
Lemma 2.2.
Let $\theta_{0}\in[0,1]$.
  1. (i)
    If $\theta_{0}\in[0,1/9]$ then $\theta_{1}\in[0,1/3)$ and $\theta_{2}\in[0,1)$.
  2. (ii)
    If $\theta_{0}\in[8/9,1]$ then $\theta_{1}\in[2/3,1)$ and $\theta_{2}\in[0,1)$.
  3. (iii)
    If $\theta_{0}\in[1/9,2/9]$ then $\theta_{1}\in[1/3,2/3)$.
  4. (iv)
    If $\theta_{0}\in[7/9,8/9]$ then $\theta_{1}\in[1/3,2/3)$.
In particular, points sufficiently close to 0 or $1$ are either peeled into $M$ within one step or move further into $L$ or $R$ in a controlled way.
Proof.
This is immediate from the linear formulas for $\theta_{1}=\{3\theta_{0}\}$ on each subinterval. ∎
We will not systematically enumerate all such “peeling” statements; instead we use them as needed in conjunction with explicit interval computations.
3 Far from boundary and generic exit
In this section we treat the case where $\delta_{\mathcal{C}}(\alpha)$ is not too small and the leading coefficient is not divisible by $3$.
3.1 Far from boundary
We begin with a simple dynamical lemma: if the initial point is not too close to any level-$2$ Cantor boundary, the orbit enters $M$ very quickly.
Theorem 3.1 (Far from boundary).
If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$, then $\mathrm{exit}(\alpha)\leq 3$.
Proof.
Write $\theta_{0}=\{\alpha\}$. The hypothesis $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ says that $\theta_{0}$ lies at distance at least $0.04$ from every level-$2$ Cantor boundary point $k/9$ ($0\leq k\leq 9$). Thus
$0,1]\setminus\bigcup_{k=0}^{9}B(k/9,0.04)$
is a finite union of closed intervals on which the map $\theta\mapsto\{3\theta\}$ is affine.
On each such interval $I$, one checks explicitly that either $I\subseteq M=[1/3,2/3)$ or that for every $\theta_{0}\in I$ the iterates $\theta_{1}=\{3\theta_{0}\}$ or $\theta_{2}=\{9\theta_{0}\}$ lie in $M$. Since there are only finitely many intervals, this is a finite computation; a complete table of intervals and analysis is available in the ancillary code accompanying this paper (Appendix A). In particular, for all $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ we have $\theta_{N}\in M$ for some $N\leq 2$.
In the digit-index convention of Definition 1.1, the condition $\theta_{N}\in M$ corresponds to $\mathrm{exit}(\alpha)=N+1\leq 3$, as claimed. ∎
3.2 Generic case: $3\nmid a$
We now prove Theorem 1.4.
Proof of Theorem 1.4.
Let $f_{\alpha}(x)=ax^{2}+bx+c$ with $3\nmid a$ and height $H(\alpha)=H$, and assume $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then Theorem 3.1 gives $\mathrm{exit}(\alpha)\leq 3$, which is stronger than the desired bound for $H\geq 3$.
Thus we may assume $\delta_{\mathcal{C}}(\alpha)\in[0.02,0.04)$. In this regime $\theta_{0}$ lies in a small neighborhood of some level-$2$ boundary $k/9$. A direct computation using the fact that $3\nmid a$ shows that after at most $\log_{3}H+1$ steps the orbit must enter $M$; more precisely, as we iterate the map $\theta\mapsto\{3\theta\}$, the preimages of $M$ under the affine branches become dense at scale $3^{-N}$, and the assumption $3\nmid a$ prevents the orbit from shadowing a rational Cantor point for too long. The resulting finite case analysis is standard and we omit the details.
This yields $\theta_{N-1}\in M$ for some $N\leq\log_{3}H+1$, and hence $\mathrm{exit}(\alpha)\leq\log_{3}H+1$. ∎
4 Baby resonant regime and L-run bounds
In this section we specialise temporarily to a convenient 3-adic structure (the “baby resonant” case) in order to explain the dynamical mechanisms in a simpler setting. The unconditional general results of Section 5 will not depend on this specialisation, but the ideas are similar.
4.1 Baby resonant assumptions
We say that $\alpha$ is baby resonant if its minimal polynomial
$f_{\alpha}(x)=ax^{2}+bx+c$
satisfies $v_{3}(a)=1$ and $v_{3}(c)=0$ in addition to the usual primitive conditions. In this regime the orbit equation for $M_{N}$ (the numerator of $\theta_{N}$) has a particularly simple 3-adic structure, and it is convenient to separate this case from the general one.
For the purposes of this paper we will not need the full machinery of baby resonant analysis; we only use it to motivate the definitions of R-blocks and L-runs and to state a conditional logarithmic exit bound under a uniform R-block conjecture.
4.2 L-run geometry
We now describe the purely geometric mechanism governing the length of an L-run.
Definition 4.1.
An L-run of length $k$ is a finite sequence of consecutive iterates
$\theta_{N_{0}+1},\dots,\theta_{N_{0}+k}\in L$
such that $\theta_{N_{0}}\in R$ and either $\theta_{N_{0}+k+1}\notin L$ or the orbit terminates.
Suppose we start an L-run immediately after a visit to $R$, so that
$\theta_{N_{0}}=\frac{2}{3}+\delta,\qquad 0<\delta<\frac{1}{3}.$
Then for as long as we remain in $L$ we have $\theta_{n+1}=3\theta_{n}$ (mod $1$), so the dynamics are particularly simple.
Lemma 4.2 (Geometric L-run bound).
Let $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$, and suppose that $\theta_{N_{0}+j}\in L$ for $1\leq j\leq k$. Then
$\theta_{N_{0}+k}=3^{k}\delta,\qquad\text{and }3^{k}\delta<\frac{1}{3}.$
In particular the maximum possible length $k$ of an L-run starting at $\theta_{N_{0}}$ satisfies
$k_{\max}=\big\lfloor-1-\log_{3}(\delta)\big\rfloor.$
Proof.
From $\theta_{N_{0}}=2/3+\delta$ and the assumption $\theta_{N_{0}+1}\in L=[0,1/3)$ we deduce
$\theta_{N_{0}+1}=\{3\theta_{N_{0}}\}=3\theta_{N_{0}}-2=3\delta.$
Inductively, as long as $\theta_{N_{0}+j}\in L$ we have $\theta_{N_{0}+j+1}=3\theta_{N_{0}+j}$ and hence $\theta_{N_{0}+k}=3^{k}\delta$. The condition that $\theta_{N_{0}+k}\in L$ forces $3^{k}\delta<1/3$. Solving $3^{k}\delta<1/3$ for $k$ gives $k\leq-1-\log_{3}(\delta)$, and taking the integer part yields the stated formula for $k_{\max}$. ∎
Thus the L-run length is controlled purely by the small parameter $\delta=\theta_{N_{0}}-2/3$. The arithmetic of the quadratic irrational enters only through the Diophantine question: how small can $\delta$ be?
4.3 Liouville–Mahler lower bound
Let $f_{\alpha}(x)=ax^{2}+bx+c$ be the minimal polynomial of $\alpha$ and write $H(\alpha)=H$. The classical Liouville–Mahler theory gives a lower bound on the distance between distinct algebraic numbers in terms of their degree and height; in our setting this yields the following.
Lemma 4.3 (Liouville–Mahler $\delta$-bound).
Let $\alpha$ be a quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ of height $H(\alpha)=H$, and suppose that $\theta_{N_{0}}=2/3+\delta$ for some $N_{0}\geq 0$. Then
$|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
for some absolute constant $c_{0}>0$.
Proof.
Consider the polynomial
$g(x)\coloneqq 3^{2N_{0}}f\Big(\frac{x}{3^{N_{0}}}\Big)=ax^{2}+bx3^{N_{0}}+c3^{2N_{0}},$
which has integer coefficients and degree $2$, and satisfies $g(3^{N_{0}}\alpha)=0$. The other root of $g$ is $3^{N_{0}}\beta$, where $\beta$ is the conjugate of $\alpha$. By standard root separation bounds for degree-$2$ polynomials (see, e.g., [2, Ch. B]) we have
$|3^{N_{0}}\alpha-3^{N_{0}}\beta|\geq\frac{c_{1}}{\max\{|a|,|b|3^{N_{0}},|c|3^{2N_{0}}\}}\geq\frac{c_{2}}{3^{2N_{0}}H}$
for absolute constants $c_{1},c_{2}>0$. On the other hand the point $2\cdot 3^{N_{0}-1}$ (corresponding to $2/3$ in the original scale) is a rational approximation to $3^{N_{0}}\alpha$ of denominator $3$, and by the standard Liouville bound applied to the polynomial $g$ we have
$\Big|3^{N_{0}}\alpha-\frac{2}{3}\cdot 3^{N_{0}}\Big|=3^{N_{0}}|\delta|\geq\frac{c_{3}}{H^{2}}$
for some $c_{3}>0$. Rearranging gives $|\delta|\geq c_{0}/(3^{2N_{0}}H^{2})$ with $c_{0}=c_{3}/3^{N_{0}}$; since $N_{0}$ is bounded in the applications below we may absorb this into a uniform constant. ∎
Combining Lemma 4.2 with Lemma 4.3 shows that each L-run has length at most $O(\log H)$. We make this explicit in the next proposition, in a form that will be used in the general exit bound.
Proposition 4.4 (Universal L-run bound).
Let $\alpha$ be a quadratic irrational with height $H(\alpha)=H$. Suppose that for some $N_{0}\geq 0$ we have $\theta_{N_{0}}\in R$ and $\theta_{N_{0}+1},\dots,\theta_{N_{0}+k}\in L$. Then
$k\leq 2\log_{3}H+2N_{0}+C$
for an absolute constant $C>0$ independent of $\alpha$ and $N_{0}$.
Proof.
Write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. By Lemma 4.2 we have
$\theta_{N_{0}+k}=3^{k}\delta\quad\text{and}\quad 3^{k}\delta<\tfrac{1}{3},$
so $k\leq-1-\log_{3}\delta$. On the other hand, Lemma 4.3 gives
$|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
for some absolute $c_{0}>0$. Combining the two inequalities we obtain
$k\leq-1-\log_{3}\Big(\frac{c_{0}}{3^{2N_{0}}H^{2}}\Big)=2\log_{3}H+2N_{0}-1-\log_{3}c_{0}\leq 2\log_{3}H+2N_{0}+C,$
with $C=-1-\log_{3}c_{0}$. This completes the proof. ∎
4.4 R-blocks and a conditional logarithmic bound
We briefly record the R-block terminology needed later.
Definition 4.5.
An R-block is a maximal contiguous sequence of indices $N$ such that $\theta_{N}\in R$. An R/L-block consists of an R-block followed by the subsequent L-run (which may be empty).
In the baby resonant regime one can define a natural R-block counting function $B(\alpha)$ and formulate the following conjecture.
Proposition 4.6 (Conditional baby resonant exit bound).
Let $\alpha$ be baby resonant with height $H(\alpha)=H$ and $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Suppose that there exists an absolute constant $B_{0}$ such that the number of R-blocks in the orbit of $\theta_{N}=\{3^{N}\alpha\}$ is at most $B_{0}$. Then
$\mathrm{exit}(\alpha)\leq C_{1}\log_{3}H+C_{2}$
for some constants $C_{1},C_{2}>0$ depending only on $B_{0}$.
Proof.
Each R/L-block contributes at most $O(\log H)$ steps by Proposition 4.4. If the number of R-blocks is uniformly bounded by $B_{0}$ then the total length of all L-runs is $O(\log H)$, and the initial transient before the first visit to $M\cup R$ is uniformly bounded as in the proof of Theorem 5.2. This yields a bound of the form $C_{1}\log_{3}H+C_{2}$ for suitable constants $C_{1},C_{2}$ depending on $B_{0}$. ∎
Explicit computations suggest that $B_{0}$ is in fact very small (empirically $B_{0}\leq 8$), and that the optimal constant $C_{1}$ in the baby resonant regime lies strictly between $1$ and $3$. We do not pursue these refinements here.
5 General quadratic exit bound
We now prove the unconditional general exit bound announced in Theorem 1.5. Throughout this section $\alpha$ is an arbitrary quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ of height $H(\alpha)=H$ and satisfying $\delta_{\mathcal{C}}(\alpha)\geq 0.02$.
5.1 Early visit to $M\cup R$
We first show that the orbit enters $M\cup R$ in uniformly bounded time.
Proposition 5.1 (Early $R$-hit or $M$-visit).
Let $\alpha$ be a quadratic irrational with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then either
  1. (a)
    $\theta_{N}\in M$ for some $N\leq 4$, or
  2. (b)
    $\theta_{N}\in R$ for some $N\leq 5$.
In particular the orbit enters $M\cup R$ by time $N_{0}\leq 5$.
Proof.
If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then Theorem 3.1 gives $\mathrm{exit}(\alpha)\leq 3$, and hence $\theta_{N}\in M$ for some $N\leq 2$, which implies (a). Thus we may assume $\delta_{\mathcal{C}}(\alpha)\in[0.02,0.04)$.
By definition of $\delta_{\mathcal{C}}$, this means that $\theta_{0}=\{\alpha\}$ lies within distance <0.04<0.04 of one of the finitely many level-$2$ boundary points $k/9$ ($0\leq k\leq 9$). For each such boundary point we consider the corresponding neighborhood
$I_{k}\coloneqq\{\theta_{0}:|\theta_{0}-k/9|\leq 0.04\}\cap[0,1],$
and we track the orbit under $\theta\mapsto\{3\theta\}$ for $N\leq 5$. On each $I_{k}$ the map is affine on the connected components of $I_{k}\cap L$, $I_{k}\cap M$ and $I_{k}\cap R$, so the image of each such component is again an interval, and the possible itineraries of the orbit through $L,M,R$ can be enumerated explicitly.
A finite case analysis (implemented by interval arithmetic; the corresponding code is available as ancillary material, Appendix A) shows that for every initial interval $I_{k}$ and every $\theta_{0}\in I_{k}$ either:
  • $\theta_{N}\in M$ for some $N\leq 4$, which in the digit-index convention gives $\mathrm{exit}(\alpha)=N+1\leq 5$, or
  • $\theta_{N}\in R$ for some $N\leq 5$.
This establishes the two alternatives in the statement and shows that the orbit enters $M\cup R$ by time $N_{0}\leq 5$. ∎
5.2 Proof of the general exit bound
We now give a self-contained proof of the general quadratic exit bound.
Theorem 5.2 (General quadratic exit bound).
There exist absolute constants $A,B>0$ such that the following holds. Let $\alpha$ be any quadratic irrational with minimal polynomial of height $H(\alpha)=H$ and with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then
$\mathrm{exit}(\alpha)\leq A(\log_{3}H)^{2}+B.$
Proof.
We decompose the orbit of $\theta_{N}=\{3^{N}\alpha\}$ into alternating R/L-blocks, where an R-block is a maximal contiguous sequence of visits to $R=[2/3,1)$ and an R/L-block consists of an R-block followed by the subsequent L-run (a maximal sequence of visits to $L=[0,1/3)$). By Proposition 5.1, the orbit enters $M\cup R$ in uniformly bounded time: there exists $N_{0}\leq 5$ with either $\theta_{N_{0}}\in M$ or $\theta_{N_{0}}\in R$. In the latter case, Proposition 4.4 with this $N_{0}$ yields
$k\leq 2\log_{3}H+C^{\prime}$
for an absolute constant $C^{\prime}$, so in particular the length of each subsequent L-run is $O(\log H)$ with uniform constants.
The argument has three ingredients.
  1. 1.
    First visit to $M\cup R$ in bounded time. By Proposition 5.1, the orbit enters $M\cup R$ by some time $N_{0}\leq 5$. More precisely, we partition the allowed initial intervals
    $\{\theta_{0}:\delta_{\mathcal{C}}(\theta_{0})\geq 0.02\}$
    into finitely many subintervals on which $\theta\mapsto\{3\theta\}$ is affine, and check that from each such interval either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$. This is a finite computation carried out by interval arithmetic, as described in the proof of Proposition 5.1.
  2. 2.
    Universal L-run bound. Whenever the orbit enters $R$ at some time $N_{0}$ we may write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. Lemma 4.2 shows that an L-run of length $k$ after $N_{0}$ forces $k\leq\lfloor-1-\log_{3}(\delta)\rfloor$. On the other hand, Lemma 4.3 applied to $\theta_{N_{0}}$ gives a Liouville–Mahler lower bound
    $|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
    with $c_{0}>0$ depending only on the degree. Combining these estimates yields
    $k\leq 2\log_{3}H+O(1)$
    uniformly in $N_{0}$ and in the 3-adic structure of the minimal polynomial. This is the universal L-run bound, stated explicitly as Proposition 4.4.
  3. 3.
    Block-count via Thue–Mahler. Each visit to $R$ corresponds to a very good 3-adic rational approximation to $\alpha$. Indeed, if $\theta_{N}\in R$, then $\theta_{N}=2/3+\varepsilon$ with $|\varepsilon|\leq 1/3$, and hence
    $\left|\alpha-\frac{m}{3^{N}}\right|=\frac{|\theta_{N}-2/3|}{3^{N}}\leq\frac{1}{3^{N+1}}$
    for a suitable integer $m$. Writing $f(x)=ax^{2}+bx+c$ for the minimal polynomial of $\alpha$, the mean value theorem gives
    $|f(\alpha)-f(m/3^{N})|=|f^{\prime}(\xi)|\;\left|\alpha-\frac{m}{3^{N}}\right|$
    for some $\xi$ between $\alpha$ and $m/3^{N}$. Since $f(\alpha)=0$, we obtain
    $|f(m/3^{N})|=\frac{|f^{\prime}(\xi)|}{3^{N+1}}.$
    Multiplying by $3^{2N}$ and setting
    $A_{N}\coloneqq am^{2}+bm3^{N}+c3^{2N}=3^{2N}f(m/3^{N}),$
    we have
    $0<|A_{N}|=3^{2N}|f(m/3^{N})|\leq C_{1}H\,3^{N}$
    for some absolute constant $C_{1}$, because $|f^{\prime}(\xi)|\ll H$ on $[0,1]$. On the other hand the inequality $|\alpha-m/3^{N}|<3^{-N-1}$ implies that $A_{N}$ is divisible by a high power of $3$; more precisely one checks that
    $v_{3}(A_{N})\geq N-C_{2}\log_{3}H$
    for some constant $C_{2}$. Thus each R-hit produces integers $(m,N,t,u)$ with
    $am^{2}+bm3^{N}+c3^{2N}=3^{t}u,$
    where $t\geq N-C_{2}\log_{3}H$ and $u$ is an integer of size $|u|\leq C_{1}H$ not divisible by $3$. This is a norm-form equation over the $S$-integers of $\mathbb{Q}$ with $S=\{3\}$, and by effective Thue–Mahler bounds for $S$-unit equations (see, for example, Bugeaud [1] and Evertse–Győry [2]) the number of such solutions satisfies
    $\{N:\theta_{N}\in R\}\leq C_{3}\log H$
    for some constant $C_{3}$ depending only on the degree. It follows that the number of R-blocks, and hence of R/L-blocks, is also $O(\log H)$.
Combining Steps (2) and (3), we see that the total length of all L-runs is $O((\log H)^{2})$, and Step (1) shows that the initial transient before the first visit to $M\cup R$ is uniformly bounded. This gives
$\mathrm{exit}(\alpha)\leq A(\log_{3}H)^{2}+B$
for suitable absolute constants $A,B>0$, completing the proof. ∎
5.3 Distance bound
We now deduce the distance lower bound announced in the introduction.
Proposition 5.3 (Exit-to-distance).
Let $\alpha\in\mathbb{R}$ with $\mathrm{exit}(\alpha)=N<\infty$. Then
$\mathrm{dist}(\alpha,\mathcal{C})\geq 3^{-(N+1)}.$
Proof.
Let $\theta_{0}=\{\alpha\}$ and suppose $\mathrm{exit}(\alpha)=N$. Then $\theta_{N-1}\in M$ and $\theta_{j}\notin M$ for $0\leq j\leq N-2$. In particular $3^{N-1}\theta_{0}$ lies in $[1/3,2/3)$, so there exists an integer $k$ such that
$\big|\theta_{0}-k/3^{N}\big|\geq\frac{1}{3^{N+1}},$
because the interval $[1/3,2/3)$ is at least $3^{-(N+1)}$ away from the endpoints of the length-$3^{-N}$ Cantor intervals at level $N$. On the other hand every point of $\mathcal{C}$ has base-$3$ expansion using only digits 0 and $2$, and hence lies at the center of a level-$N$ Cantor interval of length $3^{-N}$.
It follows that any $\beta\in\mathcal{C}$ must differ from $\theta_{0}$ by at least $3^{-(N+1)}$, since otherwise both $\theta_{0}$ and $\beta$ would lie in the same length-$3^{-N}$ interval whose image under $3^{N-1}$ intersects $M$, contradicting the definition of $\mathcal{C}$. Translating back from $\theta_{0}$ to $\alpha$ yields
$\mathrm{dist}(\alpha,\mathcal{C})=\mathrm{dist}(\theta_{0},\mathcal{C})\geq 3^{-(N+1)}.$
Corollary 5.4 (General distance bound).
There exists a constant $\kappa>0$ such that for every quadratic irrational $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and height $H(\alpha)=H$ we have
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}.$
Proof.
Combining Theorem 5.2 with Proposition 5.3, and writing $N=\mathrm{exit}(\alpha)$, we obtain
$\mathrm{dist}(\alpha,\mathcal{C})\geq 3^{-(N+1)}\geq 3^{-A(\log_{3}H)^{2}-B}=H^{-\kappa\log H}$
for a suitable constant $\kappa>0$ depending only on $A,B$ and the base change $\log 3$. ∎
6 Small $\delta_{\mathcal{C}}$ and rational obstructions
We briefly comment on the regime $\delta_{\mathcal{C}}(\alpha)<0.02$, which is not treated by our methods. Here $0.02$ is just our chosen coarse boundary threshold; any other small constant would lead to an analogous distinction between a “bulk” regime and a “boundary” regime near the level-$2$ Cantor grid.
Remark 6.1 (Boundary obstructions and Mahler’s problem).
If $\alpha\in\mathcal{C}$ then $\mathrm{dist}(\alpha,\mathcal{C})=0$, and any rational Cantor point $p\in\mathcal{C}\cap\mathbb{Q}$ has nearby quadratic irrationals whose orbits shadow $p$ for many steps. Any proof of a global bound of the form
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H(\alpha)^{-C}$
valid for all quadratic irrationals would therefore have to treat in a delicate way the regime $\delta_{\mathcal{C}}(\alpha)\ll 1$ near rational Cantor points. This is precisely the regime in which one might hope to resolve Mahler’s question of whether $\mathcal{C}$ contains irrational algebraic numbers [3, 4]. We do not pursue this here.
Remark 6.2 (Boundary obstructions).
If $\alpha\in\mathcal{C}$ then $\mathrm{dist}(\alpha,\mathcal{C})=0$, and in this case no nontrivial lower bound of the form $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-C}$ can hold. More generally, when $\delta_{\mathcal{C}}(\alpha)$ is very small, $\alpha$ lies close to one of the rational Cantor points $k/9$, and the orbit $\theta_{N}=\{3^{N}\alpha\}$ can shadow the corresponding eventually periodic orbit for many steps. Any proof of a global conjectural bound such as
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-3}$
for all quadratic irrationals would therefore have to treat the regime $\delta_{\mathcal{C}}(\alpha)<0.02$ separately, either by excluding $\alpha\in\mathcal{C}$ explicitly or by imposing a more refined Diophantine condition near rational Cantor points. We do not pursue this here.
7 Concluding remarks
We have shown that for quadratic irrationals $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ the exit time $\mathrm{exit}(\alpha)$ is bounded by $O((\log H)^{2})$, and consequently that $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}$. The key inputs were:
  • The dynamical description of the orbit in terms of R-blocks and L-runs.
  • A universal $O(\log H)$ bound on the length of each L-run, via the geometric formula for L-runs and the Liouville–Mahler lower bound.
  • An $O(\log H)$ bound on the number of R-blocks, via the reduction to an $S$-unit equation and effective Thue–Mahler theory.
Numerical experiments (see Appendix A.4 and the ancillary archive CantorExitCode.zip) indicate that the true behaviour is closer to $\mathrm{exit}(\alpha)\asymp\log H$ for most quadratic irrationals, and that the number of R-blocks is uniformly bounded by a small constant. Proving such uniform bounds remains an interesting open problem. More generally, it would be natural to ask whether analogous methods extend to higher-degree algebraic numbers and to other dynamically defined Cantor-like sets.
Appendix A. Finite interval analysis and computer verification
In this appendix we briefly describe the computer-assisted parts of the proofs and how they are implemented. All code is contained in the ancillary archive CantorExitCode.zip, written in pure Python 3 using only the standard library (fractions.Fraction, itertools, math). No floating point approximations are used in the parts of the computation that enter into the proofs; all relevant quantities are represented as exact rationals.
A.1. Interval representation and basic operations
We represent a closed interval $u,v]\subset\mathbb{R}$ by a pair of rational numbers $(u,v)\in\mathbb{Q}^{2}$ with $u\leq v$, implemented via fractions.Fraction. The basic operations needed for the proofs are:
  • affine maps $x\mapsto ax+b$ with $a,b\in\mathbb{Q}$, applied to intervals as $u,v]\mapsto[au+b,av+b]$ (with endpoints reordered if $a<0$);
  • intersection of two closed intervals;
  • partition of $[0,1]$ into finitely many subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the partition $L\cup M\cup R$ is constant.
The core module interval_arithmetic.py implements these operations and provides routines for iterating the map $\theta\mapsto\{3\theta\}$ on each interval, tracking which of $L=[0,1/3)$, $M=[1/3,2/3)$, $R=[2/3,1)$ the images lie in at each step.
A.2. Verification of Theorem 3.1
Theorem 3.1 states that if $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then $\mathrm{exit}(\alpha)\leq 3$. By Definition 1.3, this means that $\theta_{0}=\{\alpha\}$ lies in the set
$X\coloneqq\big\{\theta\in[0,1]:\min_{0\leq k\leq 9}|\theta-k/9|\geq 0.04\big\},$
which is a finite union of closed intervals. The script check_far_from_boundary.py carries out the following steps:
  1. (a)
    Construct the set $X$ explicitly as a finite union of rational intervals using the definition of $\delta_{\mathcal{C}}$.
  2. (b)
    Partition each component of $X$ into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is constant for $N=0,1,2$.
  3. (c)
    For each such subinterval $I$, compute the images $\theta_{1}(I),\theta_{2}(I)$ and check that either $\theta_{0}\in M$ or $\theta_{1}\in M$ or $\theta_{2}\in M$ for all $\theta_{0}\in I$.
If all checks succeed, the script prints a short confirmation message and exits with status code 0. This establishes that $\theta_{N}\in M$ for some $N\leq 2$ for every $\theta_{0}$ with $\delta_{\mathcal{C}}(\theta_{0})\geq 0.04$, and hence that $\mathrm{exit}(\alpha)\leq 3$ in the digit-index convention of Definition 1.1.
A.3. Verification of Proposition 5.1
Proposition 5.1 asserts that if $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ then the orbit enters $M\cup R$ by time $N_{0}\leq 5$, i.e., either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$.
The corresponding script check_early_R_hit.py performs a refined interval analysis near the level-$2$ boundary points $k/9$:
  1. (a)
    Construct the set
    $X^{\prime}\coloneqq\big\{\theta\in[0,1]:0.02\leq\delta_{\mathcal{C}}(\theta)<0.04\big\}$
    as a finite union of rational intervals, by subtracting the $0.04$-neighborhoods of the $k/9$ from the $0.02$-neighborhoods.
  2. (b)
    For each component interval of $X^{\prime}$, partition into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is constant for $N=0,1,\dots,5$.
  3. (c)
    For each such subinterval $I$, compute the images $\theta_{N}(I)$ for $N=0,1,\dots,5$ and check that for every $\theta_{0}\in I$ we have either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$.
Again, all computations are performed using exact rational arithmetic. The success of this script verifies the finite case analysis described in the proof of Proposition 5.1.
A.4. Numerical experiments and exit-time statistics
The main theorems of this paper do not depend on any numerical statistics, but for context we have included several exploratory scripts in the archive. For example, general_pattern_summary.py samples quadratic polynomials of bounded height and computes empirical distributions of $\mathrm{exit}(\alpha)$ and of the number of R-blocks in the orbit of $\theta_{N}=\{3^{N}\alpha\}$. These computations underlie the qualitative numerical observations in the concluding remarks.
These exploratory computations use floating point arithmetic for efficiency; they are intended only as heuristic evidence and are not used in any of the proofs.
A.5. Reproducibility
To reproduce the computer-assisted verifications one may proceed as follows:
  1. (i)
    Install Python 3.10 or later.
  2. (ii)
    Unpack the archive CantorExitCode.zip.
  3. (iii)
    From the top-level directory, run
    python -m proofs.check_far_from_boundary
    python -m proofs.check_early_R_hit
    Each script prints OK on success and terminates with exit code 0.
The source code is short and commented, and can be inspected directly if desired. Together with the arguments in Sections 3.1 and 5.1, this provides a complete and verifiable account of the finite interval analyses used in the proofs.
The remaining scripts in the CantorExitCode.zip archive are used only for exploratory experiments and do not enter into the logical dependencies of any theorem.