### v2

# Explicit separation of quadratic irrationals from the middle-third Cantor set

Abstract
Assuming a mild non-degeneracy condition excluding very low-level Cantor endpoints, and assuming a counting/input hypothesis for the contribution of non-deep orbit indices (Hypothesis 5.13), we show that for the quadratic field $K=\mathbb{Q}(\alpha)$ there exist constants $A_{K},B_{K}>0$ such that
$\mathrm{exit}(\alpha)\ \leq\ A_{K}\,(\log_{3}H)^{2}+B_{K}.$
Consequently, $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa_{K}\log H}$ for some $\kappa_{K}>0$.
MSC 2020: 11J04 (primary), 11D59, 28A80
Keywords: Cantor set, quadratic irrationals, Mahler’s problem
1 Introduction
Let $\mathcal{C}\subset[0,1]$ denote the middle-third Cantor set. In a 1984 problem list, Mahler asked whether $\mathcal{C}$ contains any irrational algebraic numbers and, more generally, how well algebraic numbers can approximate points of $\mathcal{C}$ [8]. A quantitative formulation is to bound $\mathrm{dist}(\alpha,\mathcal{C})$ from below in terms of the arithmetic complexity of an algebraic number $\alpha$.
In this paper we treat the case of quadratic irrationals $\alpha\in(0,1)$. Write $\theta_{0}:=\{\alpha\}$ and consider the base-$3$ dynamical system
$\tau(x)=\{3x\}\qquad(x\in[0,1]).$
Partition $[0,1]$ into the three standard intervals
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1).$
Define the exit time $\mathrm{exit}(\alpha)$ to be the least $N\geq 1$ such that $\theta_{N-1}:=\tau^{N-1}(\theta_{0})\in M$. Equivalently, $\mathrm{exit}(\alpha)$ is the position of the first ternary digit $1$ in the expansion of $\theta_{0}$.
A key observation is that $\mathrm{exit}(\alpha)$ controls $\mathrm{dist}(\alpha,\mathcal{C})$. Indeed, if $N=\mathrm{exit}(\alpha)$ then $\theta_{N-1}\in M$ while $\theta_{0},\dots,\theta_{N-2}\in L\cup R$, so $\theta_{0}$ lies in a level-(N1)(N-1) Cantor interval and enters its middle third at step $N$. This yields the exact scaling identity of Proposition 5.18, and hence lower bounds for $\mathrm{dist}(\alpha,\mathcal{C})$ once $N$ is controlled.
To obtain explicit bounds we impose a mild non-degeneracy condition:
$\delta_{\mathcal{C}}(\alpha)\ \geq\ 0.02,$
where $\delta_{\mathcal{C}}(\alpha)$ is the distance from $\theta_{0}$ to the level-$2$ Cantor endpoints (Definition 1.1). This excludes only a fixed finite union of short intervals and avoids a finite number of low-level exceptional patterns.
Definition 1.1.
Let
$E_{2}\coloneqq\{k/9:0\leq k\leq 9\},$
the set of level-$2$ Cantor boundary points in $[0,1]$. For $\alpha\in\mathbb{R}$, write $\theta_{0}=\{\alpha\}\in[0,1)$ and define
$\delta_{\mathcal{C}}(\alpha)\ \coloneqq\ \mathrm{dist}(\theta_{0},E_{2})=\min_{0\leq k\leq 9}\bigl|\{\alpha\}-k/9\bigr|.$
Main results
Let $f(x)=ax^{2}+bx+c$ be the primitive integer minimal polynomial of $\alpha$ and let $H=\max\{|a|,|b|,|c|\}$. Our main quantitative statement is an explicit polylogarithmic bound on the first time the orbit enters $M$.
Theorem 1.2 (Exit-time bound for quadratic irrationals).
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H$. Suppose that $\alpha$ has exit clearance $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Let $K=\mathbb{Q}(\alpha)$. Then there exist constants $A_{K},B_{K}>0$, depending only on $K$, such that
$\mathrm{exit}(\alpha)\ \leq\ A_{K}\,(\log_{3}H)^{2}+B_{K}.$
Assuming Hypothesis 5.13, Theorem 1.2 is proved in Theorem 5.17 below. Combining it with the exit–distance identity and a quantitative clearance bound at the exit step (Lemma 5.19) gives an explicit separation from $\mathcal{C}$.
Corollary 1.3 (A distance bound to $\mathcal{C}$).
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H(\alpha)=H$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$, and put $K=\mathbb{Q}(\alpha)$. Then there exists a constant $\kappa_{K}>0$, depending only on $K$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Idea of the proof
The dynamical obstruction to early exit is that the orbit visits the right interval $R$ and then spends a long time in $L$ before returning to $R$ (rather than entering $M$). We show that sufficiently long such “RR–then–LL” blocks force $\alpha$ to satisfy a Thue–Mahler type equation for a binary quadratic form at a 3-power argument. Effective bounds for these equations (as in [5], and algorithmically via [7]) yield unconditional control of the contribution from sufficiently deep blocks. The remaining (non-deep) contribution is isolated as Hypothesis 5.13. Assuming this hypothesis, we obtain the stated exit bound, and hence the distance bound.
2 Dynamical setup and basic lemmas
In this section we record the basic dynamical facts about the map $\theta\mapsto\{3\theta\}$ and its interaction with the partition $L\cup M\cup R$.
Definition 2.1 (Exit time).
Let $\tau\colon[0,1)\to[0,1)$ be the map $\tau(\theta)=\{3\theta\}$ and set
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1).$
For $\alpha\in\mathbb{R}$, write $\theta_{0}=\{\alpha\}$ and $\theta_{N}=\tau^{N}(\theta_{0})=\{3^{N}\alpha\}$. Define
$\mathrm{exit}(\alpha)\ \coloneqq\ \inf\{N\geq 1:\ \theta_{N-1}\in M\}\in\mathbb{N}\cup\{\infty\}.$
Remark 2.2 (Ternary digits and the ternary tree).
For $\theta\in[0,1)$ (and in particular for $\theta_{0}=\{\alpha\}$ when $\alpha$ is irrational), the map $\tau(\theta)=\{3\theta\}$ acts as the left shift on the base-$3$ expansion of $\theta$. Equivalently, $\theta$ determines an infinite path in the ternary tree whose level-$N$ node records the first $N$ ternary digits of $\theta$. The partition $L\cup M\cup R$ corresponds to the first ternary digit being $0,1,2$, respectively. In this language, the middle-third Cantor set $\mathcal{C}$ consists of paths avoiding the digit $1$ forever, and $\mathrm{exit}(\alpha)$ is the first time that the shifted orbit $\theta_{N-1}$ lands in $M$, i.e. the first occurrence of a ternary digit $1$ along the orbit itinerary.
2.1 Digit index convention
Note Definition 2.1 and the digit-index convention. For completeness we record the following simple observation.
Remark 2.3.
By definition we have $\mathrm{exit}(\alpha)=N$ if and only if $\theta_{N-1}\in M$ and $\theta_{j}\notin M$ for $0\leq j\leq N-2$. Thus bounds on the first time $N_{0}$ with $\theta_{N_{0}}\in M$ translate into exit-time bounds via $\mathrm{exit}(\alpha)=N_{0}+1$.
We will systematically work in the $\theta_{N}$-index convention and translate back to $\mathrm{exit}(\alpha)$ only at the end of proofs.
2.2 Peeling near the endpoints
We recall the (almost disjoint) decomposition
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1),$
with the understanding that boundary points are irrelevant for the orbits we consider (for quadratic irrationals, $\theta_{N}$ is irrational for all $N$).
On each of these regions the map $\tau(\theta)=\{3\theta\}$ is affine. In particular, if $\theta_{0}\in L\setminus\{1/3\}$ then $\theta_{1}=3\theta_{0}\in[0,1)$, if $\theta_{0}\in M$ then $\theta_{1}=3\theta_{0}-1$, and if $\theta_{0}\in R$ then $\theta_{1}=3\theta_{0}-2$.
We will frequently use the following simple but useful observation.
Lemma 2.4.
Let $\theta_{0}\in[0,1)$.
  1. (i)
    If $\theta_{0}\in[0,1/9]$ then $\theta_{1}\in[0,1/3)$ and $\theta_{2}\in[0,1)$.
  2. (ii)
    If $\theta_{0}\in[8/9,1)$ then $\theta_{1}\in[2/3,1)$ and $\theta_{2}\in[0,1)$.
  3. (iii)
    If $\theta_{0}\in[1/9,2/9]$ then $\theta_{1}\in[1/3,2/3)$.
  4. (iv)
    If $\theta_{0}\in[7/9,8/9]$ then $\theta_{1}\in[1/3,2/3)$.
In particular, points sufficiently close to 0 or $1$ are either peeled into $M$ within one step or move further into $L$ or $R$ in a controlled way.
Proof.
This is immediate from the linear formulas for $\theta_{1}=\{3\theta_{0}\}$ on each subinterval. ∎
We will not systematically enumerate all such “peeling” statements; instead we use them as needed in conjunction with explicit interval computations.
3 Far from boundary and generic exit
In this section we treat the case where $\delta_{\mathcal{C}}(\alpha)$ is not too small and the leading coefficient is not divisible by $3$.
3.1 Far from boundary
We begin with a simple dynamical lemma: if the initial point is not too close to any level-$2$ Cantor boundary, the orbit enters $M$ very quickly.
Theorem 3.1 (Far from boundary).
If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$, then $\mathrm{exit}(\alpha)\leq 3$.
Proof.
Let $\theta_{0}=\{\alpha\}\in[0,1)$. The condition $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ means that $\theta_{0}$ lies in the closed set
$X\coloneqq\Bigl\{\theta\in[0,1):\min_{0\leq k\leq 9}\bigl|\theta-k/9\bigr|\geq 0.04\Bigr\},$
which is a finite union of closed intervals obtained by removing open $0.04$-neighborhoods of the level-$2$ endpoints $k/9$.
On each connected component $I$ of $X$ the maps $\tau(\theta)=\{3\theta\}$ and $\tau^{2}$ restrict to affine maps, and the itinerary of points of $I$ with respect to the partition $L\cup M\cup R$ is constant. In Appendix A.2 we give a complete interval verification showing that for every such component $I$ one has
$I\subseteq M\quad\text{or}\quad\tau(I)\subseteq M\quad\text{or}\quad\tau^{2}(I)\subseteq M.$
Therefore for every $\theta_{0}\in X$ there exists $N\leq 3$ such that $\theta_{N-1}\in M$, i.e. $\mathrm{exit}(\alpha)\leq 3$. ∎
4 Baby resonant regime and L-run bounds
In this section we specialise temporarily to a convenient 3-adic structure (the “baby resonant” case) in order to explain the dynamical mechanisms in a simpler setting. The unconditional general results of Section 5 will not depend on this specialisation, but the ideas are similar.
4.1 Baby resonant assumptions
We say that $\alpha$ is baby resonant if its minimal polynomial
$f_{\alpha}(x)=ax^{2}+bx+c$
satisfies $v_{3}(a)=1$ and $v_{3}(c)=0$ in addition to the usual primitive conditions. In this regime the orbit equation for $M_{N}$ (the numerator of $\theta_{N}$) has a particularly simple 3-adic structure, and it is convenient to separate this case from the general one.
For the purposes of this paper we will not need the full machinery of baby resonant analysis; we only use it to motivate the definitions of R-blocks and L-runs and to state a conditional logarithmic exit bound under a uniform R-block conjecture.
4.2 L-run geometry
We now describe the purely geometric mechanism governing the length of an L-run.
Definition 4.1.
An L-run of length $k$ is a finite sequence of consecutive iterates
$\theta_{N_{0}+1},\dots,\theta_{N_{0}+k}\in L$
such that $\theta_{N_{0}}\in R$ and either $\theta_{N_{0}+k+1}\notin L$ or the orbit terminates.
Suppose we start an L-run immediately after a visit to $R$, so that
$\theta_{N_{0}}=\frac{2}{3}+\delta,\qquad 0<\delta<\frac{1}{3}.$
Then for as long as we remain in $L$ we have $\theta_{n+1}=3\theta_{n}$ (mod $1$), so the dynamics are particularly simple.
Lemma 4.2 (Geometric L-run bound).
Let $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$, and suppose that $\theta_{N_{0}+j}\in L$ for $1\leq j\leq k$. Then
$\theta_{N_{0}+k}=3^{k}\delta,\qquad\text{and }3^{k}\delta<\frac{1}{3}.$
In particular the maximum possible length $k$ of an L-run starting at $\theta_{N_{0}}$ satisfies
$k_{\max}=\big\lfloor-1-\log_{3}(\delta)\big\rfloor.$
Proof.
From $\theta_{N_{0}}=2/3+\delta$ and the assumption $\theta_{N_{0}+1}\in L=[0,1/3)$ we deduce
$\theta_{N_{0}+1}=\{3\theta_{N_{0}}\}=3\theta_{N_{0}}-2=3\delta.$
Inductively, as long as $\theta_{N_{0}+j}\in L$ we have $\theta_{N_{0}+j+1}=3\theta_{N_{0}+j}$ and hence $\theta_{N_{0}+k}=3^{k}\delta$. The condition that $\theta_{N_{0}+k}\in L$ forces $3^{k}\delta<1/3$. Solving $3^{k}\delta<1/3$ for $k$ gives $k\leq-1-\log_{3}(\delta)$, and taking the integer part yields the stated formula for $k_{\max}$. ∎
Thus the L-run length is controlled purely by the small parameter $\delta=\theta_{N_{0}}-2/3$. The arithmetic of the quadratic irrational enters only through the Diophantine question: how small can $\delta$ be?
4.3 Liouville–Mahler lower bound
Let $f_{\alpha}(x)=ax^{2}+bx+c$ be the minimal polynomial of $\alpha$ and write $H(\alpha)=H$. The classical Liouville–Mahler theory gives a lower bound on the distance between distinct algebraic numbers in terms of their degree and height; in our setting this yields the following.
Lemma 4.3 ($\delta$-lower bound from integrality).
Let $\alpha$ be a quadratic irrational with primitive minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H(\alpha)=H$. Let $\theta_{N}=\{3^{N}\alpha\}$ and suppose that for some $N\geq 0$ we have
$\theta_{N}=\frac{2}{3}+\delta\in R,\qquad\text{so }0<\delta<\frac{1}{3}.$
Then there is an absolute constant $c_{\delta}>0$ (independent of $\alpha$ and $N$) such that
$\delta\geq\frac{c_{\delta}}{H^{2}\,3^{N+1}}.$
Proof.
Let $p_{N},q_{N}$ be as in Definition 5.2 and set
$A_{N}\coloneqq q_{N}^{2}f(p_{N}/q_{N})=ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}.$
By Lemma 5.4, $A_{N}\in\mathbb{Z}\setminus\{0\}$ and
$|A_{N}|\leq C_{\ast}\,H^{2}\,q_{N}\,|\theta_{N}-\tfrac{2}{3}|=C_{\ast}\,H^{2}\,q_{N}\,\delta.$
Since $|A_{N}|\geq 1$ and $q_{N}=3^{N+1}$, we obtain
$\delta\geq\frac{1}{C_{\ast}\,H^{2}\,3^{N+1}}.$
Thus the conclusion holds with $c_{\delta}=1/C_{\ast}$. ∎
Combining Lemma 4.2 with Lemma 4.3 shows that an LL–run beginning after an RR–visit at time $N_{0}$ has length at most $N_{0}+O(\log H)$ (more precisely, $k\leq N_{0}+2\log_{3}H+O(1)$). We make this explicit in the next proposition, in a form that will be used in the general exit bound.
Proposition 4.4 (Universal L-run bound).
Let $\alpha$ be a quadratic irrational with height $H(\alpha)=H$. Suppose that for some $N_{0}\geq 0$ we have $\theta_{N_{0}}\in R$ and $\theta_{N_{0}+1},\dots,\theta_{N_{0}+k}\in L$. Then
$k\leq N_{0}+2\log_{3}H+C$
for an absolute constant $C>0$ independent of $\alpha$ and $N_{0}$.
Proof.
Write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. By Lemma 4.2, an $L$-run of length $k$ forces $3^{k}\delta<1/3$, hence $k\leq-1-\log_{3}\delta$.
On the other hand, Lemma 4.3 gives
$\delta\geq\frac{c_{\delta}}{H^{2}\,3^{N_{0}+1}}$
for an absolute $c_{\delta}>0$. Combining the inequalities,
$k\leq-1-\log_{3}\!\Big(\frac{c_{\delta}}{H^{2}\,3^{N_{0}+1}}\Big)=N_{0}+2\log_{3}H-\log_{3}c_{\delta},$
which yields the stated bound with $C=-\log_{3}c_{\delta}$. ∎
4.4 R-blocks and a conditional logarithmic bound
We briefly record the R-block terminology needed later.
Definition 4.5.
An R-block is a maximal contiguous sequence of indices $N$ such that $\theta_{N}\in R$. An R/L-block consists of an R-block followed by the subsequent L-run (which may be empty).
In the baby resonant regime one can define a natural R-block counting function $B(\alpha)$ and formulate the following conjecture.
Proposition 4.6 (Conditional baby resonant exit bound).
Let $\alpha$ be baby resonant with height $H(\alpha)=H$ and $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Suppose that there exists an absolute constant $B_{0}$ such that the number of R-blocks in the orbit of $\theta_{N}=\{3^{N}\alpha\}$ is at most $B_{0}$. Then
$\mathrm{exit}(\alpha)\leq C_{1}\log_{3}H+C_{2}$
for some constants $C_{1},C_{2}>0$ depending only on $B_{0}$.
Proof.
Each R/L-block contributes at most $O(\log H)$ steps by Proposition 4.4. If the number of R-blocks is uniformly bounded by $B_{0}$ then the total length of all L-runs is $O(\log H)$, and the initial transient before the first visit to $M\cup R$ is uniformly bounded as in the proof of Theorem 5.17. This yields a bound of the form $C_{1}\log_{3}H+C_{2}$ for suitable constants $C_{1},C_{2}$ depending on $B_{0}$. ∎
Explicit computations suggest that $B_{0}$ is in fact very small (empirically $B_{0}\leq 8$), and that the optimal constant $C_{1}$ in the baby resonant regime lies strictly between $1$ and $3$. We do not pursue these refinements here.
5 General quadratic exit bound
We now prove the general quadratic exit bound of Theorem 5.17. Throughout this section $\alpha$ is an arbitrary quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ of height $H(\alpha)=H$ and satisfying $\delta_{\mathcal{C}}(\alpha)\geq 0.02$.
5.1 Early visit to $M\cup R$
We first show that the orbit enters $M\cup R$ in uniformly bounded time.
Proposition 5.1 (Early $R$-hit or $M$-visit).
Let $\alpha$ be a quadratic irrational with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then either
  1. (a)
    $\theta_{N}\in M$ for some $N\leq 4$, or
  2. (b)
    $\theta_{N}\in R$ for some $N\leq 5$.
In particular the orbit enters $M\cup R$ by time $N_{0}\leq 5$.
Proof.
If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then Theorem 3.1 gives $\mathrm{exit}(\alpha)\leq 3$, and hence $\theta_{N}\in M$ for some $N\leq 2$, which implies (a). Thus we may assume $\delta_{\mathcal{C}}(\alpha)\in[0.02,0.04)$.
By definition of $\delta_{\mathcal{C}}$, this means that $\theta_{0}=\{\alpha\}$ lies within distance <0.04<0.04 of one of the finitely many level-$2$ boundary points $k/9$ ($0\leq k\leq 9$). For each such boundary point we consider the corresponding neighborhood
$I_{k}\coloneqq\{\theta_{0}:|\theta_{0}-k/9|\leq 0.04\}\cap[0,1],$
and we track the orbit under $\theta\mapsto\{3\theta\}$ for $N\leq 5$. On each $I_{k}$ the map is affine on the connected components of $I_{k}\cap L$, $I_{k}\cap M$ and $I_{k}\cap R$, so the image of each such component is again an interval, and the possible itineraries of the orbit through $L,M,R$ can be enumerated explicitly.
A finite case analysis (implemented by interval arithmetic; the corresponding code is available as ancillary material, Appendix A) shows that for every initial interval $I_{k}$ and every $\theta_{0}\in I_{k}$ either:
  • $\theta_{N}\in M$ for some $N\leq 4$, which in the digit-index convention gives $\mathrm{exit}(\alpha)=N+1\leq 5$, or
  • $\theta_{N}\in R$ for some $N\leq 5$.
This establishes the two alternatives in the statement and shows that the orbit enters $M\cup R$ by time $N_{0}\leq 5$. ∎
5.2 Triadic approximations and a Thue–Mahler reduction
In this subsection we isolate the arithmetic content of a visit to $R$ and explain how sufficiently deep visits (i.e. visits very close to $2/3$, equivalently followed by long $L$-runs) give rise to a finite family of Thue–Mahler equations.
Definition 5.2 (Digit-forced triadic rational).
Let $\alpha\in\mathbb{R}$ and $\theta_{N}=\{3^{N}\alpha\}$. For $N\geq 0$ define
$q_{N}\coloneqq 3^{N+1},\qquad p_{N}\coloneqq 3\lfloor 3^{N}\alpha\rfloor+2.$
Then $p_{N}\equiv 2\pmod{3}$, so $\gcd(p_{N},q_{N})=1$.
Lemma 5.3 (R-hit gives a forced approximation).
If $\theta_{N}\in R=[2/3,1)$, write $\theta_{N}=2/3+\varepsilon_{N}$ with $\varepsilon_{N}\in[0,1/3)$. With $p_{N},q_{N}$ as in Definition 5.2,
$\alpha-\frac{p_{N}}{q_{N}}=\frac{\theta_{N}-2/3}{3^{N}}=\frac{\varepsilon_{N}}{3^{N}},\qquad\text{hence}\qquad\Bigl|\alpha-\frac{p_{N}}{q_{N}}\Bigr|\leq\frac{1}{3^{N+1}}=\frac{1}{q_{N}}.$
Lemma 5.4 (Small value of the quadratic form).
Let $\alpha$ be a quadratic irrational with primitive minimal polynomial $f(x)=ax^{2}+bx+c$ and $H(\alpha)=H=\max(|a|,|b|,|c|)$. There exists an absolute constant $C_{\ast}>0$ such that for every $N\geq 0$, with $p_{N},q_{N}$ as above,
$A_{N}\coloneqq q_{N}^{2}f(p_{N}/q_{N})=ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}\in\mathbb{Z}\setminus\{0\}$
satisfies
$|A_{N}|\leq C_{\ast}\,H^{2}\,q_{N}\,\bigl|\theta_{N}-2/3\bigr|.$
Moreover, if $\theta_{N}\in R$ and the subsequent $L$-run has length $k\geq 1$ (i.e. $\theta_{N+1},\dots,\theta_{N+k}\in L=[0,1/3)$), then
$\bigl|\theta_{N}-2/3\bigr|<3^{-(k+1)}\qquad\text{and hence}\qquad|A_{N}|\leq C_{\ast}\,H^{2}\,3^{N-k}.$
Proof.
The integrality and nonzero-ness of $A_{N}$ are clear since $p_{N}/q_{N}\in\mathbb{Q}$ and $f(p_{N}/q_{N})\neq 0$.
For the size bound, apply the mean value theorem:
$|f(p_{N}/q_{N})|=|f(p_{N}/q_{N})-f(\alpha)|=|f^{\prime}(\xi_{N})|\,\Bigl|\alpha-\frac{p_{N}}{q_{N}}\Bigr|$
for some $\xi_{N}$ between $\alpha$ and $p_{N}/q_{N}$. For quadratic $f$, we have $f^{\prime}(x)=2ax+b$. Using the crude bound $|\alpha|\leq 1+H$ (Cauchy bound) and Lemma 5.3 (which gives $|p_{N}/q_{N}|\leq|\alpha|+1$ when $\theta_{N}\in R$), we obtain a uniform estimate $|f^{\prime}(\xi_{N})|\leq C^{\prime}H^{2}$ for an absolute constant $C^{\prime}$, hence
$|A_{N}|=q_{N}^{2}|f(p_{N}/q_{N})|\leq q_{N}^{2}\,C^{\prime}H^{2}\,\Bigl|\alpha-\frac{p_{N}}{q_{N}}\Bigr|=C^{\prime}H^{2}\,q_{N}\,\bigl|\theta_{N}-2/3\bigr|.$
This proves the first inequality with $C_{\ast}=C^{\prime}$.
If $\theta_{N}=2/3+\varepsilon$ with $\varepsilon>0$ and $\theta_{N+1},\dots,\theta_{N+k}\in L$, then Lemma 4.2 gives $\theta_{N+k}=3^{k}\varepsilon$ and the condition $\theta_{N+k}\in[0,1/3)$ yields $3^{k}\varepsilon<1/3$, i.e. $\varepsilon<3^{-(k+1)}$. Substituting $q_{N}=3^{N+1}$ gives the displayed $|A_{N}|\leq C_{\ast}H^{2}3^{N-k}$. ∎
Proposition 5.5 (Deep blocks reduce to finitely many bounded-norm $3$-power equations).
Let $\alpha$ be quadratic irrational with minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H:=\max\{|a|,|b|,|c|\}$. Fix $T>0$ and set $U_{T}:=\lceil C_{\ast}\,3^{-T}\rceil$.
Suppose $N$ is a transition time with $\theta_{N}\in\mathcal{R}$, and that the subsequent $\mathcal{L}$-run has length $k\geq N+2\log_{3}H+T$. Then
$|A_{N}|\leq U_{T}.$
Equivalently, $(p_{N},q_{N})=(p_{N},3^{N+1})$ satisfies one of finitely many Diophantine equations
$ap^{2}+bp\,3^{n}+c\,3^{2n}=u,\qquad n=N+1,\quad u\in\mathbb{Z},\ 0<|u|\leq U_{T}.$
In particular, for each fixed $u\neq 0$ this may be rewritten as the quadratic norm-form condition
$(2ap+b3^{n})^{2}-\Delta_{f}\,3^{2n}=4au,\qquad\Delta_{f}:=b^{2}-4ac,$
i.e. a Pell/norm-form type equation with the $3$-power restriction on the second variable.
Proof.
By Lemma 5.4, if the $\mathcal{L}$-run following $N$ has length $k$, then
$|A_{N}|\leq C_{\ast}\,H^{2}\,3^{N-k}.$
Assume $k\geq N+2\log_{3}H+T$. Then
$3^{N-k}\leq 3^{-2\log_{3}H-T}=H^{-2}\,3^{-T}.$
Hence
$|A_{N}|\leq C_{\ast}\,H^{2}\cdot H^{-2}\,3^{-T}=C_{\ast}\,3^{-T}\leq U_{T},$
using that $A_{N}\in\mathbb{Z}\setminus\{0\}$ and $U_{T}=\lceil C_{\ast}3^{-T}\rceil$. Finally, by definition,
$A_{N}=q_{N}^{2}f\!\left(\frac{p_{N}}{q_{N}}\right)=ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}$
with $q_{N}=3^{N+1}$. Setting $n:=N+1$ and $u:=A_{N}$ gives the displayed family of equations. The discriminant rewrite is the identity
$(2ap+bq)^{2}-\Delta_{f}q^{2}=4a(ap^{2}+bpq+cq^{2}),$
applied with $q=3^{n}$. ∎
Lemma 5.6 (Quantitative (effective) bounds for bounded-norm Thue–Mahler families).
Fix $U>0$. Let $f(X)=aX^{2}+bX+c\in\mathbb{Z}[X]$ be irreducible, of height $H$, discriminant $\Delta_{f}$, and put $K=\mathbb{Q}(\sqrt{\Delta_{f}})$. Then there exist positive constants $C_{1}=C_{1}(U,K)$ and $C_{2}=C_{2}(U,K)$, effectively computable from $U$ and $K$, such that every integer triple $(p,n,u)$ with
$ap^{2}+bp\,3^{n}+c\,3^{2n}=u,\qquad n\geq 0,\qquad|u|\leq U,$
satisfies
$h(p:3^{n})\ \leq\ C_{1}\,(\log H)^{2}+C_{2}.$
In particular, $n\leq C_{1}(\log H)^{2}+C_{2}$.
Proof.
This is a standard quantitative consequence of explicit lower bounds for linear forms in logarithms (applied after reducing the quadratic norm-form equation to an $S$-unit equation, with $S$ containing the primes dividing $6a\Delta_{f}$ and the prime $3$). One convenient reference is the quantitative theory of $S$-unit equations developed by Bugeaud–Győry [4, 3] and Evertse–Győry [5].
The key point for our purposes is the dependence of the constants: the explicit bounds involve invariants of the quadratic field $K$ (in particular its regulator and discriminant) together with the set $S$ and the heights of the coefficients, which in our situation are $\ll\log H$. Thus, for a fixed quadratic irrational $\alpha$ (hence fixed $K$), the solutions satisfy a bound of the form $C_{1}(U,K)(\log H)^{2}+C_{2}(U,K)$ as stated. We do not claim that the constants are uniform in $K$ when $K$ varies with $\Delta_{f}$. ∎
Remark 5.7 (Practical resolution of the finite Thue–Mahler family).
In applications of Proposition 5.5, the parameter set
$\mathcal{U}_{T}=\{u\in\mathbb{Z}:0<|u|\leq U_{T}\}$
is finite and (for fixed $T$) absolute. Thus, for a given quadratic polynomial $f(x)=ax^{2}+bx+c$ one obtains a finite list of Thue–Mahler instances of the form
$ap^{2}+bpq+cq^{2}=u,\qquad q=3^{n},\qquad u\in\mathcal{U}_{T},$
and Lemma 5.6 guarantees that only finitely many exponents $n$ can occur. If one wishes to make this step fully explicit (e.g. to compute the exceptional deep blocks for a particular polynomial, or to extract numerical constants), modern algorithms for efficiently solving Thue–Mahler equations can be used; see Gherga–Siksek [7].
Lemma 5.8 (Effective exponent bound for the bounded norm family).
Fix a nonzero quadratic discriminant $\Delta$ and a finite set $\mathcal{U}\subset\mathbb{Z}\setminus\{0\}$. Let $f(X)=aX^{2}+bX+c\in\mathbb{Z}[X]$ be irreducible of discriminant $\Delta_{f}=\Delta$, and put $K=\mathbb{Q}(\sqrt{\Delta})$. Then there exist positive constants $A_{\mathcal{U},K}$ and $B_{\mathcal{U},K}$, effectively computable from $\mathcal{U}$ and $K$, such that every integer triple $(p,n,u)$ with
$ap^{2}+bp\,3^{n}+c\,3^{2n}=u,\qquad n\geq 0,\qquad u\in\mathcal{U},$
satisfies
$n\ \leq\ A_{\mathcal{U},K}\,(\log H_{f})^{2}+B_{\mathcal{U},K},$
where $H_{f}$ denotes the height of $f$.
Proof.
Set $U:=\max\{|u|:u\in\mathcal{U}\}$ and apply Lemma 5.6 to $f$ and $U$. This yields
$h(p:3^{n})\ \leq\ C_{1}(U,K)\,(\log H_{f})^{2}+C_{2}(U,K),$
and in particular $n\leq C_{1}(U,K)(\log H_{f})^{2}+C_{2}(U,K)$. Absorbing $C_{1},C_{2}$ into $A_{\mathcal{U},K},B_{\mathcal{U},K}$ gives the claim. ∎
Lemma 5.9 (Few $T$-deep blocks).
Fix $T\geq 1$ and let $\alpha$ be a quadratic irrational with height $H(\alpha)=H$. Assume that the forward orbit $\{\theta_{n}\}_{n\geq 0}$ avoids $M$ for a (possibly long) time interval. Let $\mathcal{N}$ be the set of transition times
$\mathcal{N}:=\bigl\{\,N\geq 0:\ \theta_{N}\in R\ \text{ and }\ \theta_{N+1}\in L\,\bigr\}.$
For $N\in\mathcal{N}$ let $k(N)\geq 1$ be the length of the ensuing LL–run, i.e.
$k(N):=\max\bigl\{\,k\geq 1:\ \theta_{N+1},\dots,\theta_{N+k}\in L\,\bigr\}.$
Say that $N\in\mathcal{N}$ is $T$-deep if
$k(N)\ \geq\ N+2\log_{3}(H)+T.$
Then the set of $T$-deep transition times is finite, and moreover:
  1. (i)
    If $N_{0}<N_{1}<\cdots<N_{m-1}$ are the $T$-deep transition times in increasing order, then
    $N_{j+1}\ \geq\ 2N_{j}+2\log_{3}(H)+T+1\qquad(j<m-1).$
    In particular $m\ll\log\bigl(N_{m-1}+2\bigr)$.
  2. (ii)
    If, in addition, every $T$-deep transition time $N$ yields a bounded-norm identity
    $ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}=u_{N}\in\mathcal{U}_{T}$
    (as in Proposition 5.5), then then Lemma 5.8 gives a bound of the form
    $N_{m-1}\ \leq\ A_{T,K}\,(\log H)^{2}+B_{T,K},$
    for constants $A_{T,K},B_{T,K}$ effectively depending on $T$ and $K=\mathbb{Q}(\alpha)$, and hence the number of $T$-deep blocks satisfies $m\ll\log\log H$.
    Furthermore, the total time spent inside all $T$-deep blocks is $O\bigl((\log H)^{2}\bigr)$.
Proof.
Let $N\in\mathcal{N}$ and write $\theta_{N}=\tfrac{2}{3}+\delta$ with $0<\delta<\tfrac{1}{3}$. Since $\theta_{N+1}=3\delta\in L$, the LL–run length $k(N)$ is the maximal $k\geq 1$ with $3^{k}\delta<\tfrac{1}{3}$, hence $k(N)=\lfloor-1-\log_{3}\delta\rfloor$ by Lemma 4.2. If $N$ is $T$-deep then $k(N)\geq N+2\log_{3}(H)+T$, so in particular the next time $\theta$ can return from $L$ to $R$ is at least time $N+k(N)+1$, yielding
$N_{\text{next}}\ \geq\ N+k(N)+1\ \geq\ 2N+2\log_{3}(H)+T+1.$
Iterating gives (i).
For (ii), apply Proposition 5.5 at each $T$-deep transition time $N$ to obtain a bounded-norm identity with $u_{N}\in\mathcal{U}_{T}$. Writing $q_{N}=3^{N+1}=3^{n}$ with $n=N+1$, Lemma 5.8 bounds $n$ (hence $N$) by $C_{1}(\log H)^{2}+C_{2}$. Combining with the doubling inequality from (i) gives $m\ll\log\log H$. Finally, the doubling estimate implies $\sum_{j<m}N_{j}\ll N_{m-1}$, and each $T$-deep block has length $k(N_{j})\ll N_{j}+\log H$, so the total time in $T$-deep blocks is $O\bigl(N_{m-1}+\log H\cdot m\bigr)=O\bigl((\log H)^{2}\bigr)$. ∎
Lemma 5.10 (Shallow total time bound).
Let $\alpha$ be irrational and write $\theta_{n}:=\{3^{n}\alpha\}$. Assume $\mathrm{exit}(\alpha)<\infty$, so that $\theta_{n}\notin M$ for all $n<\mathrm{exit}(\alpha)$.
Let
$N_{\ast}\;:=\;\min\{n\geq 0:\ \theta_{n}\in M\cup R\},$
(with $N_{\ast}=\mathrm{exit}(\alpha)$ if $\theta_{N_{\ast}}\in M$). If $\theta_{N_{\ast}}\in R$, define the set of $R$-visits after $N_{\ast}$ by
$\mathcal{R}_{\ast}(\alpha)\;:=\;\{n:\ N_{\ast}\leq n<\mathrm{exit}(\alpha)\ \text{and}\ \theta_{n}\in R\}.$
For each $n\in\mathcal{R}_{\ast}(\alpha)$ let $\ell(n)\in\omega$ be the length of the (maximal) subsequent $L$-run:
$\ell(n)\;:=\;\max\bigl\{\ell\geq 0:\ \theta_{n+1},\dots,\theta_{n+\ell}\in L\bigr\},$
(where $\ell(n)=0$ if $\theta_{n+1}\notin L$).
Then the intervals
$I_{n}\;:=\;\{n,n+1,\dots,n+\ell(n)\}\qquad(n\in\mathcal{R}_{\ast}(\alpha))$
are pairwise disjoint and
$N_{\ast},\mathrm{exit}(\alpha)-1]\ \subseteq\ \bigcup_{n\in\mathcal{R}_{\ast}(\alpha)}I_{n}.$
Consequently,
$\mathrm{exit}(\alpha)\ \leq\ N_{\ast}\ +\ \sum_{n\in\mathcal{R}_{\ast}(\alpha)}\bigl(1+\ell(n)\bigr)\ \leq\ N_{\ast}\ +\ |\mathcal{R}_{\ast}(\alpha)|\cdot\Bigl(1+\max_{n\in\mathcal{R}_{\ast}(\alpha)}\ell(n)\Bigr).$ (5.1)
In particular, if one has bounds
$|\mathcal{R}_{\ast}(\alpha)|\ \leq\ A_{1}\log H+B_{1}\qquad\text{and}\qquad\ell(n)\ \leq\ A_{2}\log H+B_{2}\ \ (\forall n\in\mathcal{R}_{\ast}(\alpha)),$
then
$\mathrm{exit}(\alpha)\ \leq\ N_{\ast}\ +\ (A_{1}\log H+B_{1})\,(A_{2}\log H+B_{2}+1),$
i.e. $\mathrm{exit}(\alpha)=O\bigl((\log H)^{2}\bigr)$.
Proof.
If $\theta_{N_{\ast}}\in M$ then $\mathrm{exit}(\alpha)=N_{\ast}$ and we are done, so assume $\theta_{N_{\ast}}\in R$.
Fix $n\in\mathcal{R}_{\ast}(\alpha)$. By definition, $I_{n}=\{n,\dots,n+\ell(n)\}$ consists of the $R$-time $n$ together with the maximal consecutive block of subsequent times in $L$. If $m\in I_{n}$ and $m^{\prime}>m$ is the next time with $\theta_{m^{\prime}}\in R$, then necessarily $m^{\prime}\geq n+\ell(n)+1$, so the intervals $\{I_{n}\}_{n\in\mathcal{R}_{\ast}(\alpha)}$ are disjoint.
Now take any $t$ with $N_{\ast}\leq t<\mathrm{exit}(\alpha)$. Since $t<\mathrm{exit}(\alpha)$ we have $\theta_{t}\notin M$. If $\theta_{t}\in R$ then $t\in I_{t}$ (with $\ell(t)\geq 0$). If $\theta_{t}\in L$, move backwards from $t$ until leaving $L$; because $M$ is avoided before $\mathrm{exit}(\alpha)$, the first non-$L$ time must be in $R$, giving some $n\in\mathcal{R}_{\ast}(\alpha)$ with $n<t\leq n+\ell(n)$, hence $t\in I_{n}$. This proves $N_{\ast},\mathrm{exit}(\alpha)-1]\subseteq\bigcup_{n\in\mathcal{R}_{\ast}(\alpha)}I_{n}$.
Finally, disjointness implies
$\mathrm{exit}(\alpha)-N_{\ast}\ \leq\ \sum_{n\in\mathcal{R}_{\ast}(\alpha)}|I_{n}|\ =\ \sum_{n\in\mathcal{R}_{\ast}(\alpha)}(1+\ell(n)),$
which yields the stated inequalities. ∎
5.3 Proof of the general exit bound
Lemma 5.11 (Stratified Thue–Mahler reduction for transition times).
Let $\alpha$ be a quadratic irrational with primitive minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H(\alpha)=H$. Let $N\in\mathcal{N}$ be a transition time (so $\theta_{N}\in R$) with ensuing LL–run length $k(N)\geq 1$, and let
$A_{N}\coloneqq ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}\qquad\text{where}\qquad q_{N}=3^{N+1}$
as in Lemma 5.4. Define the deficiency
$d(N)\ \coloneqq\ \Bigl\lceil\,N+2\log_{3}H-k(N)\,\Bigr\rceil\in\mathbb{Z}.$
Then
$0<|A_{N}|\ \leq\ C_{\ast}\,3^{d(N)},$
where $C_{\ast}$ is the constant from Lemma 5.4. Consequently, for each fixed integer $d$, every transition time $N$ with $d(N)=d$ gives a solution $(p,n)=(p_{N},N+1)$ of one of the finitely many Thue–Mahler equations
$ap^{2}+bp3^{n}+c3^{2n}=u,\qquad n\geq 1,\qquad u\in\mathbb{Z},\ 0<|u|\leq\bigl\lceil C_{\ast}3^{d}\bigr\rceil,$
i.e. a finite union of Thue–Mahler families with $S=\{3\}$.
Proof.
Lemma 5.4 gives $0<|A_{N}|\leq C_{\ast}H^{2}3^{N-k(N)}$. Since $H^{2}=3^{2\log_{3}H}$, this is
$|A_{N}|\leq C_{\ast}\,3^{\,N+2\log_{3}H-k(N)}.$
By definition of $d(N)=\lceil N+2\log_{3}H-k(N)\rceil$ we have $3^{\,N+2\log_{3}H-k(N)}\leq 3^{d(N)}$, hence $|A_{N}|\leq C_{\ast}3^{d(N)}$. The final claim is immediate since $A_{N}\in\mathbb{Z}\setminus\{0\}$. ∎
Lemma 5.12 (Quadratic norm form $\Rightarrow$ $S$-unit equation).
Let $\alpha$ be quadratic with primitive minimal polynomial $f(x)=ax^{2}+bx+c$ and discriminant $\Delta_{f}=b^{2}-4ac\neq 0$. Put $K:=\mathbb{Q}(\sqrt{\Delta_{f}})$ and let $(\cdot)^{\prime}$ denote the nontrivial $K/\mathbb{Q}$ conjugation. Set $\omega:=\sqrt{\Delta_{f}}\in K$ and $\beta:=b+\omega$.
Let $S$ be the set of places of $K$ consisting of all archimedean places and all finite primes of $K$ above the rational primes dividing $6a\Delta_{f}$. Let $\mathcal{O}_{K,S}$ be the ring of $S$-integers and $\mathcal{O}_{K,S}^{\ast}$ its group of $S$-units.
Fix a finite set $U\subset\mathbb{Z}\setminus\{0\}$. Then there exists a finite set $\Gamma(U)\subset\mathcal{O}_{K,S}\setminus\{0\}$ with the following property: for every triple $(p,n,u)$ with $p\in\mathbb{Z}$, $n\geq 1$, $u\in U$ satisfying
$ap^{2}+bp\,3^{n}+c\,3^{2n}\;=\;u,$ (5.2)
there exist $\gamma\in\Gamma(U)$ and $\varepsilon\in\mathcal{O}_{K,S}^{\ast}$ such that
$z\;:=\;2ap+\beta\,3^{n}\;=\;\gamma\,\varepsilon.$ (5.3)
Moreover, defining
$\lambda:=\frac{\gamma^{\prime}}{\gamma}\in K^{\times},\qquad\mu:=\frac{2\omega}{\gamma}\in K^{\times},\qquad x:=\frac{\varepsilon^{\prime}}{\varepsilon}\in\mathcal{O}_{K,S}^{\ast},\qquad y:=\frac{3^{n}}{\varepsilon}\in\mathcal{O}_{K,S}^{\ast},$
we obtain the $S$-unit equation
$\lambda\,x+\mu\,y\;=\;1.$ (5.4)
Proof.
Multiplying (5.2) by $4a$ and completing the square gives the standard identity
$(2ap+b3^{n})^{2}-\Delta_{f}\,3^{2n}\;=\;4au.$
In $K=\mathbb{Q}(\omega)$ this factors as
$(2ap+(b+\omega)3^{n})\,(2ap+(b-\omega)3^{n})\;=\;4au,$
i.e. $z\,z^{\prime}=4au$ with $z:=2ap+\beta\,3^{n}$. By the definition of $S$ (which in particular inverts the primes above $2$, 33, the primes dividing $a$, and the primes dividing $\Delta_{f}$), we have $z\in\mathcal{O}_{K,S}$. Hence the principal ideal $(z)\subset\mathcal{O}_{K,S}$ divides the ideal $(4au)$ in $\mathcal{O}_{K,S}$.
Because $U$ is finite, the set of ideals dividing $(4au)$ (for $u\in U$) is finite. Choose once and for all a generator $\gamma\in\mathcal{O}_{K,S}$ for each such principal ideal; let $\Gamma(U)$ be the resulting finite set of chosen generators. Then for any solution of (5.2) with $u\in U$ we have $(z)=(\gamma)$ for some $\gamma\in\Gamma(U)$, hence $z=\gamma\varepsilon$ for some $\varepsilon\in\mathcal{O}_{K,S}^{\ast}$, proving (5.3).
Finally, subtracting the conjugate relation $z^{\prime}=\gamma^{\prime}\varepsilon^{\prime}$ from $z=\gamma\varepsilon$ gives $z-z^{\prime}=2\omega\,3^{n}=\gamma\varepsilon-\gamma^{\prime}\varepsilon^{\prime}$. Dividing by $\gamma\varepsilon$ yields
$1-\frac{\gamma^{\prime}}{\gamma}\cdot\frac{\varepsilon^{\prime}}{\varepsilon}\;=\;\frac{2\omega}{\gamma}\cdot\frac{3^{n}}{\varepsilon},$
which is exactly (5.4) with the displayed definitions of $\lambda,\mu,x,y$. ∎
Hypothesis 5.13 (Shallow contribution bound).
Fix $T\geq 1$. Let $\alpha$ be a quadratic irrational of height $H(\alpha)=H$ satisfying $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Let $N_{*}$, $\mathcal{R}_{*}(\alpha)$, and $\ell(n)$ be as in Lemma 5.10.
Define the set of $T$-deep indices
$\mathcal{R}_{\mathrm{deep}}:=\{n\in\mathcal{R}_{*}(\alpha):\ \ell(n)\geq n+2\log_{3}H+T\},$
and let $\mathcal{R}_{\mathrm{sh}}:=\mathcal{R}_{*}(\alpha)\setminus\mathcal{R}_{\mathrm{deep}}$.
Then the total contribution from the non-$T$-deep indices before exit satisfies
$\sum_{n\in\mathcal{R}_{\mathrm{sh}}}\bigl(1+\ell(n)\bigr)=O\bigl((\log H)^{2}\bigr),$
with an implied constant depending only on $T$ and the quadratic field $\mathbb{Q}(\alpha)$.
Remark 5.14.
Hypothesis 5.13 is the only additional input used in Step 3 of the proof of Theorem 5.17. Our methods give unconditional control of the $T$-deep contribution (via reduction to Thue–Mahler/S-unit equations and quantitative finiteness), but do not currently provide a matching unconditional bound for the total contribution of the remaining (non-deep) indices.
Remark 5.15 (Unconditional content).
For each fixed $T\geq 1$, the analysis of $T$-deep transitions is unconditional: Lemma 5.9 bounds the number of $T$-deep transitions by $O(\log\log H)$ and their total time contribution by $O_{K}((\log H)^{2})$.
Lemma 5.16 (Quantitative input for two-term SS–unit equations).
Let $K$ be a number field and $S$ a finite set of places of $K$ containing all archimedean places. Let $\lambda,\mu\in K^{\times}$ and consider the two-term SS–unit equation
$\lambda x+\mu y\;=\;1,\qquad x,y\in\mathcal{O}_{K,S}^{\ast}.$ (5.5)
Then:
  1. (i)
    (5.5) has only finitely many solutions $(x,y)\in(\mathcal{O}_{K,S}^{\ast})^{2}$. Moreover, the number of solutions admits an explicit upper bound depending only on $K:\mathbb{Q}]$ and $|S|$ (hence effectively on $K$ and $S$).
  2. (ii)
    There are explicit upper bounds for the heights $h(x)$ and $h(y)$ of all solutions in terms of $K,S$ and the coefficient heights $h(\lambda),h(\mu)$ (equivalently, in terms of $K,S$ and $h(\lambda:\mu:1)$).
In particular, in the quadratic situation of Lemma 5.12 (where $K=\mathbb{Q}(\sqrt{\Delta_{f}})$ and $S$ is fixed), these bounds yield effective control on the valuations at primes above $3$ of any SS–unit solution $y$, and hence on exponents $n$ occurring when $y$ is of the form $y=3^{n}/\varepsilon$ with $\varepsilon\in\mathcal{O}_{K,S}^{\ast}$.
Proof.
This is standard in the theory of SS–unit equations. Finiteness and explicit bounds on the number of solutions follow from quantitative forms of the Subspace Theorem (for fixed $K,S$), while effective height bounds for the solutions follow from Baker-type lower bounds for linear forms in logarithms specialized to SS–units. For explicit treatments in the present level of generality (including effective constants), see Bugeaud–Győry [4] and Evertse–Győry [5]. The final assertion follows by expressing $y$ in a basis of fundamental SS–units and reading off its 33–adic valuation from the exponent vector, which is bounded in terms of the solution height. ∎
We now give a self-contained proof of the general quadratic exit bound.
Theorem 5.17 (Log-square exit bound).
Assume Hypothesis 5.13. Let $K$ be a quadratic number field. Then there exist constants $A_{K},B_{K}>0$, depending only on $K$, such that for every $\alpha\in K\cap(0,1)$ with primitive minimal polynomial $f$ and $\delta_{\mathcal{C}}(\alpha)\geq 0.02$,
$\mathrm{exit}(\alpha)\ \leq\ A_{K}(\log_{3}H(f))^{2}+B_{K}.$
Proof.
Let $\theta_{0}=\{\alpha\}\in[0,1)$ and $\theta_{n}=\{3^{n}\alpha\}$. By Proposition 5.1, there exists $N_{0}\leq 5$ with $\theta_{N_{0}}\in M\cup R$. If $\theta_{N_{0}}\in M$ then $\mathrm{exit}(\alpha)\leq N_{0}+1$ and we are done. Thus assume $\theta_{N_{0}}\in R$.
We decompose the time before exit using Lemma 5.10. Let $N_{*}$, $\mathcal{R}_{*}(\alpha)$, and $\ell(n)$ be as there. Let
$\mathcal{N}:=\{\,N\in\mathcal{R}_{*}(\alpha):\ \ell(N)\geq 1\,\}$
be the set of transition times from $R$ into $L$, and for $N\in\mathcal{N}$ write $k(N):=\ell(N)$.
Step 1: a universal run-length bound. By Proposition 4.4 we have
$k(N)\ \leq\ N+2\log_{3}(H)+C_{0}$ (5.6)
for an absolute constant $C_{0}$.
Step 2: deep transitions and Thue–Mahler. Fix an integer parameter $T\geq 1$ to be chosen later, and call $N\in\mathcal{N}$ $T$-deep if
$k(N)\ \geq\ N+2\log_{3}H+T.$
For such an $N$, Proposition 5.5 yields a bounded-norm identity
$ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}=u_{N},\qquad q_{N}=3^{N+1},\qquad u_{N}\in\mathcal{U}_{T}.$
Let
$\mathcal{U}_{T}\ :=\ \{u\in\mathbb{Z}:\ 0<|u|\leq U_{T}\}.$
Applying Lemma 5.8 (with the fixed finite set $\mathcal{U}_{T}$ from Proposition 5.5) yields an upper bound
$N\;\leq\;A_{T,\mathbb{Q}(\alpha)}(\log H)^{2}+B_{T,\mathbb{Q}(\alpha)},$
where the constants are effective and depend only on the deepness threshold $T$ and the quadratic field $\mathbb{Q}(\alpha)$ (equivalently on the quadratic field determined by the squarefree part of $\Delta_{f}$).
Moreover, Lemma 5.9 implies that the number of $T$-deep transitions is $m\ll\log\log H$, and the total time spent inside all $T$-deep $R/L$-blocks is
$\text{(deep block time)}\ =\ O\bigl((\log H)^{2}\bigr).$ (5.7)
Step 3: the remaining (non-deep) indices. Let
$\mathcal{R}_{\mathrm{deep}}:=\{n\in\mathcal{R}_{*}(\alpha):\ \ell(n)\geq n+2\log_{3}H+T\},\qquad\mathcal{R}_{\mathrm{sh}}:=\mathcal{R}_{*}(\alpha)\setminus\mathcal{R}_{\mathrm{deep}}.$
By Hypothesis 5.13,
$\sum_{n\in\mathcal{R}_{\mathrm{sh}}}\bigl(1+\ell(n)\bigr)=O\bigl((\log H)^{2}\bigr).$
Together with (5.7) and (5.1), this completes the proof of the exit bound.
Conclusion. Starting from (5.1) and splitting the sum into deep and non-deep indices, we combine (5.7) with Hypothesis 5.13 to obtain
$\mathrm{exit}(\alpha)\ \leq\ A_{T,K}(\log_{3}H)^{2}+B_{T,K}.$
Now fix once and for all a deepness threshold $T=T_{0}$ (e.g. $T_{0}=1$), and set $A_{K}:=A_{T_{0},K}$ and $B_{K}:=B_{T_{0},K}$, which depend only on $K=\mathbb{Q}(\alpha)$. This completes the proof. ∎
5.4 Distance bound
We now deduce the distance lower bound announced in the introduction.
Proposition 5.18 (Exit-to-distance: exact scaling).
Let $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ and suppose $\mathrm{exit}(\alpha)=N<\infty$. Write $\theta_{n}=\{3^{n}\alpha\}\in[0,1)$. Then $\theta_{N-1}\in M=[1/3,2/3)$ and
$\mathrm{dist}(\alpha,\mathcal{C})=\mathrm{dist}(\theta_{0},\mathcal{C})=\frac{1}{3^{N-1}}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr).$
In particular, $\mathrm{dist}(\alpha,\mathcal{C})<3^{-N}$, and for any $\eta>0$, if $\mathrm{dist}(\theta_{N-1},\{1/3,2/3\})\geq\eta$ then
$\mathrm{dist}(\alpha,\mathcal{C})\geq\eta\,3^{-(N-1)}.$
Proof.
Set $\theta_{0}=\{\alpha\}$ and let $N=\mathrm{exit}(\alpha)$, so that $\theta_{0},\dots,\theta_{N-2}\notin M$ while $\theta_{N-1}\in M$. Since $\alpha\notin\mathbb{Q}$, all $\theta_{n}$ are irrational and hence avoid the triadic rationals; in particular each iterate lies in exactly one of $L$, $M$, or $R$.
The condition $\theta_{0},\dots,\theta_{N-2}\notin M$ means that the first N1N-1 ternary digits of $\theta_{0}$ lie in $\{0,2\}$, so $\theta_{0}$ lies in a unique level-(N1)(N-1) Cantor interval $I$ of length $3^{-(N-1)}$ on which $\tau^{N-1}$ is affine:
$\tau^{N-1}(x)=3^{N-1}x-k\qquad(x\in I)$
for some integer $k$. Moreover $\theta_{N-1}=\tau^{N-1}(\theta_{0})\in M$ says exactly that $\theta_{0}$ lies in the open middle third of the interval $I$ removed at level $N$.
The closest points of the closed set $\mathcal{C}\cap I$ to a point in this removed open middle third are the two endpoints of that middle third. These two endpoints are precisely the unique points $x_{1/3},x_{2/3}\in I$ with
$\tau^{N-1}(x_{1/3})=1/3,\qquad\tau^{N-1}(x_{2/3})=2/3.$
By the affine formula, $|\,\theta_{0}-x_{1/3}\,|=|\,\theta_{N-1}-1/3\,|/3^{N-1}$ and similarly for $2/3$. Therefore
$\mathrm{dist}(\theta_{0},\mathcal{C})=\min\bigl(|\theta_{0}-x_{1/3}|,\ |\theta_{0}-x_{2/3}|\bigr)=\frac{1}{3^{N-1}}\mathrm{dist}\!\bigl(\theta_{N-1},\{1/3,2/3\}\bigr).$
The displayed inequalities follow immediately since $\mathrm{dist}(\theta_{N-1},\{1/3,2/3\})\in(0,1/3)$. ∎
Lemma 5.19 (Exit clearance lower bound).
Let $\alpha\in(0,1)$ be a quadratic irrational with primitive integer minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H=\max\{|a|,|b|,|c|\}\geq 1$. Let $N=\mathrm{exit}(\alpha)$ and put $\theta_{N-1}:=\{3^{N-1}\alpha\}\in M=[1/3,2/3)$. Then
$\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\ \geq\ \frac{1}{30\,H^{2}\,3^{N}}.$
Proof.
Let $m=\lfloor 3^{N-1}\alpha\rfloor\in\mathbb{Z}$ and set $q:=3^{N}$. If $\theta_{N-1}\leq 1/2$, set $p:=3m+1$; otherwise set $p:=3m+2$. Then $p\in\mathbb{Z}$ and
$\left|\alpha-\frac{p}{q}\right|=\frac{1}{3^{N-1}}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr).$
Since $\alpha$ is irrational, $f(p/q)\neq 0$, so
$A:=q^{2}f(p/q)=ap^{2}+bpq+cq^{2}$
is a nonzero integer; hence $|f(p/q)|\geq q^{-2}$.
Let $\xi$ lie between $\alpha$ and $p/q$. By the mean value theorem,
$|f(p/q)|=|f(p/q)-f(\alpha)|=|f^{\prime}(\xi)|\left|\alpha-\frac{p}{q}\right|.$
Using the Cauchy bound $|\alpha|\leq 1+H$ (and $p/q\in(0,1)$), we have $|\xi|\leq 2+H$ and therefore
$|f^{\prime}(\xi)|=|2a\xi+b|\leq 2H(2+H)+H\leq 5H^{2}.$
Consequently
$\left|\alpha-\frac{p}{q}\right|\geq\frac{1}{5H^{2}q^{2}}.$
Multiplying by $3^{N-1}$ yields
$\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\geq 3^{N-1}\cdot\frac{1}{5H^{2}\cdot 3^{2N}}=\frac{1}{15H^{2}\cdot 3^{N}},$
and the displayed bound with $30$ follows by weakening constants. ∎
Corollary 5.20 (Distance bound in terms of $H$ alone).
Let $\alpha\in(0,1)$ be a quadratic irrational of height $H\geq 2$ and let $N=\mathrm{exit}(\alpha)$. Then
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ \frac{1}{30\,H^{2}}\,3^{-2N}.$
In particular, if $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and assuming Hypothesis 5.13 (so that Theorem 5.17 applies), there exists a constant $\kappa_{K}>0$, depending only on $K=\mathbb{Q}(\alpha)$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Proof.
Combine Proposition 5.18 with Lemma 5.19:
$\mathrm{dist}(\alpha,\mathcal{C})=3^{-(N-1)}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\geq 3^{-(N-1)}\cdot\frac{1}{30H^{2}3^{N}}=\frac{1}{30H^{2}}\,3^{-2N+1}.$
Weakening the constant gives the displayed $\frac{1}{30H^{2}}3^{-2N}$ bound.
Now suppose $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and Hypothesis 5.13 holds, and put $K=\mathbb{Q}(\alpha)$. Apply Theorem 5.17 to bound $N\leq A_{K}(\log_{3}H)^{2}+B_{K}$, and rewrite $3^{-2A_{K}(\log_{3}H)^{2}}=H^{-(2A_{K}/\log 3)\log H}$. Absorb the remaining polynomial factor $H^{-2}$ and constants into a single exponent $\kappa_{K}$. ∎
Corollary 5.21 (Distance to $\mathcal{C}$ from the exit-time bound).
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H(\alpha)=H$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$, and put $K=\mathbb{Q}(\alpha)$. Then there exists a constant $\kappa_{K}>0$, depending only on $K$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Proof.
This is exactly the second conclusion of Corollary 5.20. ∎
6 Small $\delta_{\mathcal{C}}$ and rational obstructions
We briefly comment on the regime $\delta_{\mathcal{C}}(\alpha)<0.02$, which is not treated by our methods. Here $0.02$ is just our chosen coarse boundary threshold; any other small constant would lead to an analogous distinction between a “bulk” regime and a “boundary” regime near the level-$2$ Cantor grid.
Remark 6.1 (Boundary obstructions and Mahler’s problem).
If $\alpha\in\mathcal{C}$ then $\mathrm{dist}(\alpha,\mathcal{C})=0$, and any rational Cantor point $p\in\mathcal{C}\cap\mathbb{Q}$ has nearby quadratic irrationals whose orbits shadow $p$ for many steps. Any proof of a global bound of the form
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H(\alpha)^{-C}$
valid for all quadratic irrationals would therefore have to treat in a delicate way the regime $\delta_{\mathcal{C}}(\alpha)\ll 1$ near rational Cantor points. This is precisely the regime in which one might hope to resolve Mahler’s question of whether $\mathcal{C}$ contains irrational algebraic numbers [8, 9]. We do not pursue this here.
Remark 6.2 (Boundary obstructions).
If $\alpha\in\mathcal{C}$ then $\mathrm{dist}(\alpha,\mathcal{C})=0$, and in this case no nontrivial lower bound of the form $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-C}$ can hold. More generally, when $\delta_{\mathcal{C}}(\alpha)$ is very small, $\alpha$ lies close to one of the rational Cantor points $k/9$, and the orbit $\theta_{N}=\{3^{N}\alpha\}$ can shadow the corresponding eventually periodic orbit for many steps. Any proof of a global conjectural bound such as
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-3}$
for all quadratic irrationals would therefore have to treat the regime $\delta_{\mathcal{C}}(\alpha)<0.02$ separately, either by excluding $\alpha\in\mathcal{C}$ explicitly or by imposing a more refined Diophantine condition near rational Cantor points. We do not pursue this here.
Further directions: game-theoretic largeness on Cantor-type sets
A complementary line of work studies Diophantine properties on fractals using Schmidt-type games and their variants adapted to Cantor constructions. In particular, Badziahin–Harrap introduced the framework of Cantor-winning sets, which enjoy robustness properties analogous to classical winning sets (e.g. stability under countable intersections and bi-Lipschitz changes of variables) while being tailored to Cantor-like splitting structures [1]. More recently, Badziahin–Harrap–Nesharim–Simmons developed a systematic comparison between Schmidt games and Cantor-winning notions [2]. These game-theoretic notions address robust largeness inside fractals, whereas the present paper is concerned with quantitative separation for a countable family (quadratic irrationals); nevertheless, the shared “ternary tree” viewpoint suggests a useful parallel language for future work near the boundary regime.
7 Concluding remarks
Conditional on Hypothesis 5.13, we have shown that for quadratic irrationals $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ the exit time $\mathrm{exit}(\alpha)$ is bounded by $O((\log H)^{2})$, and consequently that $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa_{K}\log H}$. The key inputs were:
  • The dynamical description of the orbit in terms of R-blocks and L-runs.
  • A universal $O(\log H)$ bound on the length of each L-run, via the geometric formula for L-runs and the Liouville–Mahler lower bound.
  • An unconditional analysis of sufficiently deep R/L-blocks via reduction to an $S$-unit equation and effective Thue–Mahler theory, together with the shallow-input Hypothesis 5.13.
Numerical experiments (see Appendix A.4 and the ancillary archive CantorExitCode.zip) indicate that the typical behaviour is closer to $\mathrm{exit}(\alpha)\asymp\log H$ in the sampled ranges, and that the number of R-blocks appears to remain small. Proving such uniform bounds, and in particular establishing Hypothesis 5.13 unconditionally, remains an interesting open problem. More generally, it would be natural to ask whether analogous methods extend to higher-degree algebraic numbers and to other dynamically defined Cantor-like sets.
Appendix A. Finite interval analysis and computer verification
In this appendix we briefly describe the computer-assisted parts of the proofs and how they are implemented. All code is contained in the ancillary archive CantorExitCode.zip, written in pure Python 3 using only the standard library (fractions.Fraction, itertools, math). No floating point approximations are used in the parts of the computation that enter into the proofs; all relevant quantities are represented as exact rationals.
The purpose of this appendix is solely to document the finite interval verifications used in the proofs (of Theorem 3.1 and Proposition 5.1). No large-scale computation is used to support any theorem, and we do not aim for exhaustive empirical exploration beyond the qualitative experiments mentioned in A.4.
A.1. Interval representation and basic operations
We represent a closed interval $u,v]\subset\mathbb{R}$ by a pair of rational numbers $(u,v)\in\mathbb{Q}^{2}$ with $u\leq v$, implemented via fractions.Fraction. The basic operations needed for the proofs are:
  • affine maps $x\mapsto ax+b$ with $a,b\in\mathbb{Q}$, applied to intervals as $u,v]\mapsto[au+b,av+b]$ (with endpoints reordered if $a<0$);
  • intersection of two closed intervals;
  • partition of $[0,1]$ into finitely many subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the partition $L\cup M\cup R$ is constant.
The core module interval_arithmetic.py implements these operations and provides routines for iterating the map $\theta\mapsto\{3\theta\}$ on each interval, tracking which of $L=[0,1/3)$, $M=[1/3,2/3)$, $R=[2/3,1)$ the images lie in at each step.
A.2. Verification of Theorem 3.1
Theorem 3.1 states that if $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then $\mathrm{exit}(\alpha)\leq 3$. By Definition 1.1, this means that $\theta_{0}=\{\alpha\}$ lies in the set
$X\coloneqq\big\{\theta\in[0,1]:\min_{0\leq k\leq 9}|\theta-k/9|\geq 0.04\big\},$
which is a finite union of closed intervals. The script proofs/check_far_from_boundary.py (run as python -m proofs.check_far_from_boundary) carries out the following steps:
  1. (a)
    Construct the set $X$ explicitly as a finite union of rational intervals using the definition of $\delta_{\mathcal{C}}$.
  2. (b)
    Partition each component of $X$ into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is constant for $N=0,1,2$.
  3. (c)
    For each such subinterval $I$, compute the images $\theta_{1}(I),\theta_{2}(I)$ and check that either $\theta_{0}\in M$ or $\theta_{1}\in M$ or $\theta_{2}\in M$ for all $\theta_{0}\in I$.
If all checks succeed, the script prints a short confirmation message and exits with status code 0. This establishes that $\theta_{N}\in M$ for some $N\leq 2$ for every $\theta_{0}$ with $\delta_{\mathcal{C}}(\theta_{0})\geq 0.04$, and hence that $\mathrm{exit}(\alpha)\leq 3$ in the digit-index convention of Definition 2.1.
A.3. Verification of Proposition 5.1
Proposition 5.1 asserts that if $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ then the orbit enters $M\cup R$ by time $N_{0}\leq 5$, i.e., either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$.
The corresponding script proofs/check_early_R_hit.py (run as python -m proofs.check_early_R_hit) performs a refined interval analysis near the level-$2$ boundary points $k/9$:
  1. (a)
    Construct the set
    $X^{\prime}\coloneqq\big\{\theta\in[0,1]:0.02\leq\delta_{\mathcal{C}}(\theta)<0.04\big\}$
    as a finite union of rational intervals, by subtracting the $0.04$-neighborhoods of the $k/9$ from the $0.02$-neighborhoods.
  2. (b)
    For each component interval of $X^{\prime}$, partition into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is constant for $N=0,1,\dots,5$.
  3. (c)
    For each such subinterval $I$, compute the images $\theta_{N}(I)$ for $N=0,1,\dots,5$ and check that for every $\theta_{0}\in I$ we have either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$.
Again, all computations are performed using exact rational arithmetic. The success of this script verifies the finite case analysis described in the proof of Proposition 5.1.
A.4. Numerical experiments and exit-time statistics
The main theorems of this paper do not depend on any numerical statistics, but for context we have included several exploratory scripts in the archive. For example, general_pattern_summary.py samples quadratic polynomials of bounded height and computes empirical distributions of $\mathrm{exit}(\alpha)$ and of the number of R-blocks in the orbit of $\theta_{N}=\{3^{N}\alpha\}$. These computations underlie the qualitative numerical observations in the concluding remarks.
These exploratory computations use floating point arithmetic for efficiency; they are intended only as heuristic evidence and are not used in any of the proofs. The scripts are parameterized (e.g. sample size and coefficient/height cutoffs); we refer the reader to the header comments in the code for the exact ranges used in our runs.
A.5. Reproducibility
To reproduce the computer-assisted verifications one may proceed as follows:
  1. (i)
    Install Python 3.10 or later.
  2. (ii)
    Unpack the archive CantorExitCode.zip.
  3. (iii)
    From the top-level directory, run
    python -m proofs.check_far_from_boundary
    python -m proofs.check_early_R_hit
    Each script prints OK on success and terminates with exit code 0.
The source code is short and commented, and can be inspected directly if desired. Together with the proofs of Theorem 3.1 and Proposition 5.1, this provides a complete and verifiable account of the finite interval analyses used in the proofs.
The remaining scripts in the CantorExitCode.zip archive are used only for exploratory experiments and do not enter into the logical dependencies of any theorem.